### v1

# Compton scattering from superstrings

Abstract
We propose a candidate Compton amplitude which is valid for any (integer) quantum spin and free from any spurious poles. We consider the cases of electromagnetism and gravity. We obtain such amplitudes by calculating the corresponding ones from superstring theory involving states on the leading Regge trajectory. To extract the associated field-theory amplitudes a few considerations in the form of simple physical constraints are required, such as: Soft momentum transfer, compactification of polarizations and consistent factorization in the physical channels. We believe the present exploration will be significantly relevant for the physics of compact binary systems with spin.
1 Introduction
An important insight arising from studies of quantum scattering amplitudes is that gravitational amplitudes are simpler than one would expect [1]. The double copy prescription determines that the gravitational amplitudes are obtained as a product of two Yang-Mills amplitudes [2, 3, 4]. The original formulation by Kawai, Lewellen and Tye states that a closed string amplitude at the tree level is given by a sum over the products of two open string amplitudes at tree level [5]. The current trend is invaluable for the physics of gravitational waves [6, 7, 8, 9] due to the birth of gravitational wave astronomy, driven by landmark observations made by the LIGO and Virgo collaborations [10, 11, 12, 13, 14].
In this context, we have witnessed a particularly intense flurry of activities about the gravitational interaction of spinning particles [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60]  [^4] For a much larger body of older and current research please see the references within [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60].. Indeed, gravitational dynamics of spinning objects has been a subject of vigorous study for several years [61, 62, 63, 64]. An outstanding problem emerging in this regard is the evaluation of the gravitational Compton scattering amplitude – the opposite-helicity version is a key ingredient in the evaluation of the conservative potential of compact binary systems with spin [41, 42, 43, 65, 66] and also waveform calculations with spin [60, 59, 67]. However, previous calculations in the literature displayed the presence of non-physical poles for massive particles with quantum spins greater than one, in the electromagnetic case, and greater than two, in the gravitational case [68, 69]  [^5] More precisely, the opposite-helicity Compton amplitude built from BCFW recursion relations have unphysical poles because this procedure is no longer predictive in the case of higher spins. This is due to the fact that the product of BCFW shifted three-point amplitudes does not approach zero fast enough as the shift parameter goes to infinity [25]. Nevertheless, we remark that if one starts with a local Lagrangian the amplitudes calculated from it using standard Feynman rules are expected to be manifestly local; see Ref. [41] for a calculation for the case of massive spin 3 electromagnetic Compton.. Important ongoing research is being conducted to address this problem [28, 25, 21, 22, 27, 29, 30, 31, 32, 33, 58].
This scenario opens up a great opportunity to investigate string amplitudes. It is known that one of the most notable features of string theory is the presence of an infinite tower of massive spinning modes in its spectrum required for consistency [70, 71]. However, some difficulties have been highlighted in the literature [72, 73, 42]; in particular, the classical limit of 3-point string amplitudes associated with leading Regge states of the open and closed superstring differ from the expected result for a Kerr black hole. This seems to indicate that using string theory to describe Kerr black holes is at the very least a formidable and challenging task.
In this paper we wish to report significant progress. As an unexpected result we extract from a string-theory based calculation the correct tree-level Compton amplitude for a Kerr black hole for quantum spin two – the analogous result for $\sqrt{\textrm{Kerr}}$ is for quantum spin one, which we also recover here  [^6] Paolo di Vecchia and Francesco Alessio have also obtained this result through a slightly different calculation. We thank them for sharing their notes with us.. Having succeeded in recovering this important benchmark, we propose a candidate Compton amplitude which is free from any spurious poles and valid to any integer quantum spin. In order to obtain such amplitudes we impose three plausible physical constraints, to be discussed later on in this paper. With these simple rules we manage to obtain sensible field-theory Compton amplitudes which are physically sound. We are employing natural units, with $\hbar=c=1$.
2 Compton scattering for open and closed strings
2.1 Open strings
Let us calculate the Compton amplitude involving two photons and two massive spin-$S$ particles. For simplicity in this paper we will only study the NS sector of the string. We consider the Minkowski metric as $\eta_{\mu\nu}=\textrm{diag}(-,+,+,\cdots,+)$ and Mandelstam variables defined as $s_{ij}=-(k_{i}+k_{j})^{2}$, with $s_{12}=s$, $s_{13}=t$ and $s_{14}=u$. In addition, $k_{1}^{2}=k_{3}^{2}=-m^{2}$. Here $s+t+u=2m^{2}=2(S-1)/\alpha^{\prime}$, and the leading Regge state satisfies $m^{2}\alpha^{\prime}=n=S-1$. The vertex operators required here are the following [71]
$V_{n}^{(-1)}(k;z)=\frac{e}{(2\alpha^{\prime})^{n/2}}\epsilon_{\mu_{1}\ldots\mu_{n}\nu}(k):\left(\prod_{j=1}^{n}i\partial X^{\mu_{j}}\right)\psi^{\nu}(z)e^{-\phi(z)}e^{ik\cdot X(z)}:$ (1)
and
$V_{0}^{(0)}(q;z)=e\sqrt{2\alpha^{\prime}}\epsilon_{\nu}(q):\left((q\cdot\psi)\psi^{\nu}(z)+\frac{1}{2\alpha^{\prime}}i\partial X^{\nu}\right)e^{iq\cdot X(z)}:.$ (2)
We employ a given representation of the massive polarization tensors $\epsilon_{(S)}$ (they are symmetric, traceless and transverse in all indices) of spin-S particles in which they are written as symmetric tensor products of massive spin-$1$ polarization vectors $\epsilon_{1}$, namely $\epsilon_{(S)}=\bigotimes_{S}\epsilon_{1}$. We should also consider $V_{0}^{(0)}$ – in order to cancel the background ghost charge on a genus $g$ Riemann surface, a $g$-loop superstring amplitude needs an overall superghost charge of $22g-$ 2. Finally, we will take $V_{0}^{(0)}$ to represent the photon operator in the zero picture. Our disk amplitude on the leading Regge trajectory reads
[…]
where the factor $e^{-\lambda}$ comes from the Euler-number term in the action. For convenience we write the amplitude as
$A_{4}({\bf 1},2^{h_{2}},{\bf 3},4^{h_{4}})$ $=$ $\epsilon_{\alpha}(2)\epsilon_{\beta}(4)\,\epsilon({\bf 3})\cdot A^{\alpha\beta}_{4}({\bf 1},2^{h_{2}},{\bf 3},4^{h_{4}})\cdot\epsilon({\bf 1})$ (4)
$=$ $\epsilon_{\alpha}(2)\epsilon_{\beta}(4)\left(A^{\alpha\beta}_{4}({\bf 1},2^{h_{2}},{\bf 3},4^{h_{4}})\right)^{a_{1}\ldots a_{n}a_{S}}\,_{c_{1}\ldots c_{n}c_{S}}\,\,\epsilon_{a_{1}\ldots a_{n}a_{S}}({\bf 1})\epsilon^{c_{1}\ldots c_{n}c_{S}}({\bf 3})$
where the dot products refer to the contraction of the higher-spin indices for legs 1 and 3. Henceforth we are suppressing the higher-spin indices for clarity. Moreover, we consider $z_{1}>z_{2}>z_{4}$. This means that we integrate $z_{3}$ over $[z_{4},z_{2}]$ to ensure cyclic ordering $(1,2,3,4)$ of the insertion points. Here we take $z_{1}\to\infty$, $z_{2}=1$ and $z_{4}=0$, which implies that $z_{3}\in[0,1]$.
We proceed to calculate the integrals by employing well known correlators in the literature [71, 74]. There are associated constants of proportionality that are fixed by requiring that the amplitude describes the interaction of photons with massive particles. Hence a straightforward calculation produces the following color-ordered amplitude
$A^{\alpha\beta}_{4}({\bf 1},2^{h_{2}},{\bf 3},4^{h_{4}})\equiv A^{\alpha\beta}_{t}({\bf 1},2^{h_{2}},{\bf 3},4^{h_{4}})=\frac{{\cal V}_{t}}{\alpha^{\prime}t\,\prod_{k=1}^{2n-1}(\alpha^{\prime}t-k)}{\cal A}_{0}^{\alpha\beta},$ (5)
where ${\cal V}_{t}\equiv\Gamma(1-s\alpha^{\prime})\Gamma(1-u\alpha^{\prime})/\Gamma(1+\alpha^{\prime}t-2n)$. The explicit expression for ${\cal A}_{0}^{\alpha\beta}$ (which does not contain propagators) is rather involved and we refer the reader to the appendix for details, as well as for some explanation concerning notation. It is easy to see that the full amplitude is invariant under the simultaneous exchange of $\alpha\leftrightarrow\beta$ and $q_{2}\leftrightarrow q_{4}$.
2.2 Closed strings
In order to display the Compton amplitude for the closed string, we will resort to the known KLT relations [75] ($N$ is the number of external particles)
$M_{N}$ $=$ $\left(\frac{-i}{4}\right)^{N-3}\sum_{\sigma}\sum_{\gamma,\beta}{\cal S}_{\alpha^{\prime}}[\gamma\bigl{(}\sigma(2),\ldots,\sigma(j-1)\bigr{)}|\sigma(2,\ldots,j-1)]_{k_{1}}$ (6)
$\times$ ${\cal S}_{\alpha^{\prime}}[\beta\bigl{(}\sigma(j),\ldots,\sigma(N-2)\bigr{)}|\sigma(j,\ldots,N-2)]_{k_{N-1}}A_{N}(1,\sigma(2,\ldots,N-2),N-1,N)\bigg{|}_{\alpha^{\prime}\to\alpha^{\prime}/4}$
$\times$ $\tilde{A}_{N}\left(\gamma\bigl{(}\sigma(2),\ldots,\sigma(j-1)\bigr{)},1,N-1,\beta\bigl{(}\sigma(j),\ldots,\sigma(n-2)\bigr{)},N\right)\bigg{|}_{\alpha^{\prime}\to\alpha^{\prime}/4}$
where we sum over all orderings, $\sigma(1,\ldots,N)$, $\gamma(1,\ldots,N)$ and $\beta(1,\ldots,N)$ denote permutations of the $N!$ labels and the gauge amplitudes $A,\tilde{A}$ are associated with the left-moving and right-moving sectors, respectively. The momentum kernel ${\cal S}_{\alpha^{\prime}}$ is defined in Ref. [75], but we quote its expression in the appendix. The amplitude can also be written as
$M_{N}$ $=$ $(-1)^{N-3}\sum_{\sigma,\gamma}{\cal S}_{\alpha^{\prime}}[\gamma(2,\ldots,N-2)|\sigma(2,\ldots,N-2)]_{k_{1}}$ (7)
$\times$ $A_{N}(1,\sigma(2,\ldots,N-2),N-1,N)\bigg{|}_{\alpha^{\prime}\to\alpha^{\prime}/4}\tilde{A}_{N}(N-1,N,\gamma(2,\ldots,N-2),1)\bigg{|}_{\alpha^{\prime}\to\alpha^{\prime}/4}.$
Recall that, on the leading Regge trajectory for the closed string, the spin is $S=2(n+1)$, where $m^{2}\alpha^{\prime}=4n$ for the mass level $n$.
For the Compton amplitude, the KLT relation reads
$M_{4}({\bf 1}^{S},2^{h_{2}\tilde{h}_{2}},{\bf 3}^{S},4^{h_{4}\tilde{h}_{4}})$ $=$ $(-1)^{n}\left(\frac{\pi\alpha^{\prime}}{2}\right)^{-1}\sin\left(\frac{\pi\alpha^{\prime}s}{4}\right)$ (8)
$\times$ $A_{4}({\bf 1}^{S/2},2^{h_{2}},{\bf 3}^{S/2},4^{h_{4}})\bigg{|}_{\alpha^{\prime}\to\alpha^{\prime}/4}\tilde{A}_{4}({\bf 1}^{S/2},2^{\tilde{h}_{2}},4^{\tilde{h}_{4}},{\bf 3}^{S/2})\bigg{|}_{\alpha^{\prime}\to\alpha^{\prime}/4}$
where we used reflection symmetry for the right-moving sector and the gauge amplitudes are color-ordered amplitudes, see appendix for definitions. In order to describe interactions with gravitons, we have to correlate the helicities in the gauge amplitudes, $h_{2}=\tilde{h}_{2}$ and $h_{4}=\tilde{h}_{4}$. Notice that we are also correlating the spins of the massive particles. In our previous notation, we can write the amplitude as
$M^{\alpha\gamma,\beta\delta}_{4}({\bf 1},2,{\bf 3},4)$ $=$ $(-1)^{n}\left(\frac{\pi\alpha^{\prime}}{2}\right)^{-1}\sin\left(\frac{\pi\alpha^{\prime}s}{4}\right)\sin\left(\frac{\pi\alpha^{\prime}u}{4}\right)$ (9)
$\times$ $\left\{\frac{{\cal V}_{t}}{\sin(\alpha^{\prime}\pi t)\alpha^{\prime}t\,\prod_{k=1}^{2n-1}(\alpha^{\prime}t-k)}\frac{{\cal V}_{t}}{\alpha^{\prime}t\,\prod_{l=1}^{2n-1}(\alpha^{\prime}t-l)}{\cal A}_{0}^{\alpha\beta}{\cal A}_{0}^{\gamma\delta}\right\}_{\alpha^{\prime}\to\alpha^{\prime}/4}$
where we used the cyclic property for the gauge color-ordered amplitude and monodromy relations. As before higher-spin indices are left implicit. The above open-string amplitudes were stripped from the factor $2ie^{2}$.
3 Electromagnetic Compton amplitudes for all spins
Our aim is to extract from the above string results terms potentially relevant for the classical description of spinning compact binary systems. Superstring amplitudes are famously defined in ten dimensions, so we need to perform a compactification down to four dimensions; we follow the prescription given by Ref. [72] and define suitable four-dimensional polarizations and momenta. Moreover, we also isolate the poles corresponding to the mass level $n$; these are connected to the s-channel and the u-channel. To extract the associated field-theory amplitudes for generic spins, we consider both momenta associated with the external photons as being soft, $q\to\hbar q$. Observe that keeping only the level $n$ poles corresponds to the leading 1/$1/\hbar$ term in the soft expansion. Then at leading order in a soft expansion we find
$A^{\alpha\beta}_{t}\Bigg{|}_{\textrm{field theory}}\equiv{\cal A}^{\alpha\beta}_{S}=\frac{1}{[\Gamma(n)]^{2}}\frac{{\cal A}_{0}^{\alpha\beta}}{(s\alpha^{\prime}-n)(u\alpha^{\prime}-n)}+{\cal O}(\alpha^{\prime}t).$ (10)
where the subscript $S$ in ${\cal A}^{\alpha\beta}_{S}$ denotes the spin of the massive particle. Before continuing the discussion for generic spins, let us study the lowest-spin massive boson. The polarization tensors are given by $\epsilon_{a_{1}a_{S}}({\bf 1})=\epsilon_{a_{1}}({\bf 1})\epsilon_{a_{S}}({\bf 1})$ and $\epsilon_{c_{1}c_{S}}({\bf 3})=\epsilon_{c_{1}}({\bf 3})\epsilon_{c_{S}}({\bf 3})$. By taking a close look at the corresponding expression, one quickly finds terms that describe the Compton amplitude for scattering of massive spin $1$ particles. To separate such terms from the rest we consider a suitable compactification of the polarization vectors $\epsilon_{a_{1}}({\bf 1})$ and $\epsilon_{c_{1}}({\bf 3})$ – we take these to be in the “fifth dimension” (such as $\epsilon^{d=10}({\bf 1})=(0,0,0,0,1,{\bf 0})$). After performing this compactification, contraction with the polarizations, and a few algebraic manipulations, we arrive at the following expression of the Compton amplitude describing the interaction between photons and massive spin-$1$ particles:
${\cal A}_{1}({\bf 1},2^{h_{2}},{\bf 3},4^{h_{4}})$ $=$ $-\frac{2ie^{2}}{(s-m^{2})(u-m^{2})}\biggl{\{}2\epsilon({\bf 1})\cdot\epsilon({\bf 3})\,k_{1}\cdot F_{2}\cdot F_{4}\cdot k_{1}$ (11)
$+$ $\left[k_{1}\cdot F_{2}\cdot k_{3}\,F_{4}^{\alpha\beta}+k_{1}\cdot F_{4}\cdot k_{3}\,F_{2}^{\alpha\beta}+\frac{1}{2}k_{1}\cdot(q_{4}-q_{2})\bigl{(}F_{2}^{\alpha\rho}F_{4\rho}^{\ \ \beta}-F_{2}^{\beta\rho}F_{4\rho}^{\ \ \alpha}\bigr{)}\right]{\cal J}_{\alpha\beta}$
-- $\frac{q_{2}\cdot q_{4}}{8}\bigl{(}F_{2\alpha\beta}F_{4\lambda\kappa}+F_{4\alpha\beta}F_{2\lambda\kappa}\bigr{)}\{{\cal J}^{\alpha\beta},{\cal J}^{\lambda\kappa}\}\biggr{\}}$
where $h_{2},h_{4}$ are the helicities of the photons and $F_{i}^{\mu\nu}=2\epsilon^{[\mu}(i)q_{i}^{\nu]}$. To obtain this expression we wrote $\epsilon_{a_{S}}({\bf 1})\epsilon_{c_{S}}({\bf 3})={\cal J}_{a_{S}c_{S}}/2+\epsilon_{(a_{S}}({\bf 1})\epsilon_{c_{S})}({\bf 3})$, where we have defined ${\cal J}^{\alpha\gamma}\equiv-i\epsilon^{\mu}({\bf 1})(M^{\alpha\gamma})_{\mu\nu}\epsilon^{\nu}({\bf 3})$, $(M^{ab})_{c}^{\ d}=2i\delta^{[a}_{\ c}\eta^{b]d}$ being the Lorentz generator in the spin-$1$ representation. Furthermore:
$\frac{1}{2}\{{\cal J}^{\alpha\beta},{\cal J}^{\lambda\kappa}\}\equiv-\epsilon^{\mu}({\bf 1})\bigl{[}(M^{\alpha\beta})_{\mu\rho}(M^{\lambda\kappa})^{\rho}_{\ \nu}+(M^{\lambda\kappa})_{\mu\rho}(M^{\alpha\beta})^{\rho}_{\ \nu}\bigr{]}\epsilon^{\nu}({\bf 3}).$ (12)
Up to a numerical overall factor, the result (11) perfectly matches the one previously derived by others [22] (except for possible sign differences due to different conventions on momenta). By rewriting this expression with the help of spinor-helicity methods, we also find that this amplitude agrees with other known expressions in the literature [68, 69].
Observe that $\alpha^{\prime}$ is effectively parameterizing the expansion in the soft photon momenta. Nevertheless, to see that the approach of taking directly $\alpha^{\prime}\to 0$ is not, in general, the proper way to obtain the field theory limit which is useful in the present context, let us quote the amplitude for the spin-2 leading Regge state. The calculation is rather involved but straightforward. After the dust settles, we find that
${\cal A}_{2}({\bf 1},2^{h_{2}},{\bf 3},4^{h_{4}})$ $=$ $-\frac{2ie^{2}}{\left(s-m^{2}\right)\left(u-m^{2}\right)}\Bigg{\{}2\epsilon({\bf 1})\cdot\epsilon({\bf 3})k_{1}\cdot F_{2}\cdot F_{4}\cdot k_{1}$ (13)
$+$ $2\Bigg{[}k_{1}\cdot F_{2}\cdot k_{3}F_{4}^{\mu\nu}+k_{1}\cdot F_{4}\cdot k_{3}F_{2}^{\mu\nu}+\frac{1}{2}k_{1}\cdot\left(q_{4}-q_{2}\right)\left(F_{2}^{\mu\rho}F_{4\rho}^{\ \ \nu}-F_{2}^{\nu\rho}F_{4\rho}^{\ \ \mu}\right)\Bigg{]}{\cal J}_{\mu\nu}$
$+$ $\frac{q_{2}\cdot q_{4}}{2}\Big{(}F_{2\mu\nu}F_{4\rho\sigma}+F_{2\rho\sigma}F_{4\mu\nu}\Big{)}\left(\frac{1}{2}\{{\cal J}^{\mu\nu},{\cal J}^{\rho\sigma}\}+8\epsilon^{\mu\sigma}({\bf 1})\epsilon^{\nu\rho}({\bf 3})\right)\Bigg{\}}$
where ${\cal J}^{ab}\equiv-\frac{i}{2}\epsilon^{c_{1}c_{2}}({\bf 1}){(M^{ab})_{c_{1}c_{2}}}^{d_{1}d_{2}}\epsilon_{d_{1}d_{2}}({\bf 3})$, ${(M^{ab})_{c(2)}}^{d(2)}=4i\delta_{(c_{1}}^{[a}\eta^{b](d_{1}}\delta_{c_{2})}^{d_{2})}$ being the Lorentz generator in the spin-$2$ representation. This amplitude in the classical limit will produce contributions up to the quadrupole term. However, for a spin-$S$ quantum particle, the classical counterpart should produce an expansion up to the $2S$-th power of the classical spin. Besides, strictly speaking, the $\alpha^{\prime}\to 0$ amounts to projecting out all massive spinning particles from the spectrum. These are clearly indications that the limit $\alpha^{\prime}\to 0$ is not adequate to our purposes.
Given the precedent successful computation, Eq. (11), let us proceed to the general case. Following previous prescriptions, we consider the compactification of two massive polarization vectors, namely $\epsilon_{a_{n}}({\bf 1})$ and $\epsilon_{c_{n}}({\bf 3})$. Furthermore, we consider only terms with $k=l$ in the double summations in the full expression for ${\cal A}_{0}^{\alpha\beta}$ (see appendix) as these are the ones that have explicit inverse propagators in the numerators. Our candidate electromagnetic Compton amplitude takes the form
${\cal A}_{S}^{\alpha\beta}$ $=$ $\frac{2ie^{2}}{\left(s-m^{2}\right)\left(u-m^{2}\right)}\sum_{j=0}^{n-1}\sum_{k=0}^{n-1-j}\frac{(-1)^{n-1-j}n^{2}}{\Gamma(j+2)\Gamma(k+1)^{2}\Gamma(-j-k+n)^{2}}$ (14)
$\times$ $(2\alpha^{\prime})^{n-1-j}\eta^{a(j)c(j)}\gamma(s\alpha^{\prime},n-1)\gamma(u\alpha^{\prime},n-1)q_{2}^{c(n-1-j-k)}q_{4}^{a(k)}q_{2}^{a(n-1-j-k)}q_{4}^{c(k)}$
$\times$ $\Biggl{\{}-(u-m^{2})\left[\Bigl{(}-\eta^{a_{s}c_{s}}k_{3}^{\beta}+\eta^{a_{s}\beta}q_{4}^{c_{s}}-\eta^{c_{s}\beta}q_{4}^{a_{s}}\Bigr{)}\Bigl{(}k_{3}^{\alpha}+q_{4}^{\alpha}\Bigr{)}+q_{2}^{a_{s}}\left(\eta^{\alpha\beta}q_{4}^{c_{s}}-\eta^{c_{s}\beta}q_{4}^{\alpha}\right)\right.$
$+$ $\left.\eta^{a_{s}\alpha}\Bigl{(}(q_{2}\cdot q_{4})\eta^{c_{s}\beta}-q_{2}^{\beta}q_{4}^{c_{s}}\Bigr{)}+k_{3}^{\beta}\left(q_{2}^{c_{s}}\eta^{a_{s}\alpha}-q_{2}^{a_{s}}\eta^{\alpha c_{s}}\right)\right]$
-- $(s-m^{2})\left[\Bigl{(}-\eta^{a_{s}c_{s}}k_{3}^{\alpha}+q_{2}^{c_{s}}\eta^{a_{s}\alpha}-q_{2}^{a_{s}}\eta^{\alpha c_{s}}\Bigr{)}\Bigl{(}k_{3}^{\beta}+q_{2}^{\beta}\Bigr{)}+q_{4}^{a_{s}}\left(\eta^{\alpha\beta}q_{2}^{c_{s}}-\eta^{\alpha c_{s}}q_{2}^{\beta}\right)\right.$
$+$ $\left.\eta^{a_{s}\beta}\Bigl{(}(q_{2}\cdot q_{4})\eta^{\alpha c_{s}}-q_{2}^{c_{s}}q_{4}^{\alpha}\Bigr{)}+k_{3}^{\alpha}\left(q_{4}^{c_{s}}\eta^{a_{s}\beta}-q_{4}^{a_{s}}\eta^{c_{s}\beta}\right)\right]$
$+$ $\frac{1}{2}\eta^{\alpha\beta}\eta^{a_{s}c_{s}}(s-m^{2})(u-m^{2})\Biggr{\}},$
where $\gamma(x,N)\equiv\prod_{m=1}^{N}(-x+m)=(1-x)_{N}$. This amplitude is gauge invariant in the sense that $q_{4,\beta}{\cal A}^{\alpha\beta}_{n}=0=q_{2,\alpha}{\cal A}^{\alpha\beta}_{n}$. The number of the mass level $n$ characterizes the spin of the massive particle under consideration and we can just take $n\to S$ in the above expressions.
There is still one more constraint that we should consider – we require the associated residues to be given by the product of two minimally coupled 3-point amplitudes. For this purpose, let us analyze the s-channel cut (similar considerations apply to the u-channel cut). We take particle 3 to be a small boost of particle 1 (or the inverse) [72]:
[…]
which implies that
$\frac{1}{m}\langle{\bf 1}{\bf 3}\rangle$ $=$ $\mathbbm{1}+\frac{1}{2m_{q}^{2}}\,\langle{\bf 1}|q|{\bf 1}\bigl{]}$
$\frac{1}{m}\bigl{[}{\bf 1}{\bf 3}\bigr{]}$ $=$ $\mathbbm{1}-\frac{1}{2m_{q}^{2}}\,\langle{\bf 1}|q|{\bf 1}\bigl{]}$
$\epsilon^{\mu}(\mathbf{3})$ $=$ $\bar{\epsilon}^{\mu}(\mathbf{1})-\frac{q\cdot\bar{\epsilon}(\mathbf{1})}{m_{q}^{2}}\left(k_{1}^{\mu}+\frac{1}{2}q^{\mu}\right)$ (16)
where $m_{q}=m\sqrt{1+q^{2}/4m^{2}}$ and $q\equiv-k_{3}-k_{1}=q_{2}+q_{4}$ is the momentum transferred, which is very small compared to the massive momenta. Therefore we can write that
$(q_{2}\cdot\epsilon({\bf 3})\,q_{2}\cdot\epsilon({\bf 1}))^{l}$$\approx$ $(q_{2}\cdot\bar{\epsilon}(\mathbf{1})\,q_{2}\cdot\epsilon({\bf 1}))^{l}$
$(q_{4}\cdot\epsilon({\bf 3})\,q_{4}\cdot\epsilon({\bf 1}))^{l}$$\approx$ $(q_{4}\cdot\epsilon({\bf 3})\,q_{4}\cdot\bar{\epsilon}(\mathbf{3}))^{l}$
$\bigl{(}\epsilon(\mathbf{3})\cdot\epsilon(\mathbf{1})\bigr{)}^{j}$ $=$ $\left(\frac{2}{m}\langle{\bf 3}{\bf 1}\rangle\right)^{j}\sum_{p=0}^{j}\frac{j!}{p!(j-p)!}\left(\frac{1}{2m}\langle{\bf 3}{\bf 1}\rangle\right)^{p}.$ (17)
The s-channel residue reads (after contracting with external polarizations)
${\cal A}_{n}\Bigg{|}_{\textrm{s-channel residue}}$$\approx$ $-2ie^{2}\sum_{j=0}^{n-1}\sum_{k=0}^{n-1-j}\sum_{p=0}^{j}\frac{j!}{p!(j-p)!}\frac{(-1)^{n-1-j}n^{2}(2n)^{n-1-j}[\gamma(n,n-1)]^{2}}{\Gamma(j+2)\Gamma(k+1)^{2}\Gamma(-j-k+n)^{2}}$ (18)
$\times$ $\left(\frac{2}{m}\langle{\bf 3}{\bf 1}\rangle\right)^{j}\left(\frac{1}{2m}\langle{\bf 3}{\bf 1}\rangle\right)^{p}\left(\frac{\bigl{(}q_{2}\cdot\bar{\epsilon}({\bf 1})\,q_{2}\cdot\epsilon({\bf 1})\bigr{)}}{m^{2}}\right)^{n-1-j-k}\left(\frac{\bigl{(}q_{4}\cdot\bar{\epsilon}({\bf 3})\,q_{4}\cdot\epsilon({\bf 3})\bigr{)}}{m^{2}}\right)^{k}$
$\times$ $\epsilon_{a_{s}}({\bf 1})\Big{[}\Bigl{(}-\eta^{a_{s}c_{s}}k_{3}\cdot\epsilon(4)+\epsilon^{a_{s}}(4)q_{4}^{c_{s}}-\epsilon^{c_{s}}(4)q_{4}^{a_{s}}\Bigr{)}(k_{3}+q_{4})\cdot\epsilon(2)$
$+$ $q_{2}^{a_{s}}\Big{(}\epsilon(2)\cdot\epsilon(4)\,q_{4}^{c_{s}}-\epsilon^{c_{s}}(4)\,q_{4}\cdot\epsilon(2)\Big{)}$
$+$ $\epsilon^{a_{s}}(2)\Bigl{(}(q_{2}\cdot q_{4})\,\epsilon^{c_{s}}(4)-q_{2}\cdot\epsilon(4)\,q_{4}^{c_{s}}\Bigr{)}$
$+$ $(k_{3}\cdot\epsilon(4))\Big{(}q_{2}^{c_{s}}\epsilon^{a_{s}}(2)-q_{2}^{a_{s}}\epsilon^{c_{s}}(2)\Big{)}\Big{]}\epsilon_{c_{s}}({\bf 3}).$
where, after using s-channel kinematics and the fact that $q^{2}\lll 1/\alpha^{\prime}$, we find that $\gamma(s\alpha^{\prime},n-1)=\gamma(n,n-1)$ and $\gamma(u\alpha^{\prime},n-1)=\gamma(n-t\alpha^{\prime},n-1)\approx\gamma(n,n-1)$ at leading order. To understand how we should impose consistent factorization as a constraint, let us analyze the corresponding “helicity amplitudes” separately. First let us consider the all plus amplitude. The term in square brackets is just the spin multipole decomposition of the QED Compton amplitude for matter of spin one [22], so this must be proportional to $\bigl{[}24\bigr{]}^{2}\langle{\bf 1}{\bf 3}\rangle^{2}/t$ [69]. So the rest of the amplitude should produce something like $m^{2-2S}\langle{\bf 1}{\bf 3}\rangle^{2S-2}$. The only term that furnishes this is given by the contribution $p=j=n-1$:
$\bar{\cal A}_{n}(++)\Bigg{|}_{\textrm{s-channel residue}}$ $=$ $2ie^{2}m^{2-2n}\bigl{[}24\bigr{]}^{2}\langle{\bf 1}{\bf 3}\rangle^{2n}\frac{1}{u-m^{2}}\left(\frac{n^{2}[\gamma(n,n-1)]^{2}}{\Gamma(n+1)}\right)$ (19)
where we have denoted by a bar such a contribution. We also used s-channel kinematics. Apart from the positive numerical term inside the brackets, this term reproduces the correct s-channel factorization. Now let us focus on the amplitude with opposite helicities. As the term in square brackets in Eq. (18) is roughly the QED Compton amplitude for matter of spin one, this should be proportional to
$\frac{\langle 4|{\bf 1}|2\bigr{]}^{2}}{m^{4}t}\big{|}\langle{\bf 1}|P|{\bf 3}\bigr{]}\big{|}^{2}$
where $P=-k_{1}-q_{2}=k_{3}+q_{4}$. However, since $|{\bf 1}\rangle$ and $|\bar{\bf 1}\rangle$ are related by complex conjugation (the same for ${\bf 3}$), we find that
$q_{2}\cdot\bar{\epsilon}({\bf 1})\,q_{2}\cdot\epsilon({\bf 1})$ $=$ $-\frac{|\langle{\bf 1}|P|{\bf 1}\bigr{]}|^{2}}{2m^{2}}\approx-\frac{|\langle{\bf 1}|P|{\bf 3}\bigr{]}|^{2}}{2m^{2}}$
$q_{4}\cdot\bar{\epsilon}({\bf 3})\,q_{4}\cdot\epsilon({\bf 3})$ $=$ $-\frac{|\langle{\bf 3}|P|{\bf 3}\bigr{]}|^{2}}{2m^{2}}\approx-\frac{|\langle{\bf 1}|P|{\bf 3}\bigr{]}|^{2}}{2m^{2}},$ (20)
up to subleading terms in the classical limit, where we used that $k_{1}\cdot\epsilon({\bf 1})=k_{1}\cdot\bar{\epsilon}({\bf 1})=0$, and similarly for ${\bf 3}$. Only the term $j=0$ reproduces the proper s-channel residue; hence
$\bar{\cal A}_{n}(+-)\Bigg{|}_{\textrm{s-channel residue}}\approx 2ie^{2}\frac{\langle 4|{\bf 1}|2\bigr{]}^{2}}{m^{4n}(u-m^{2})}\big{|}\langle{\bf 1}|P|{\bf 3}\bigr{]}\big{|}^{2n}\left(\frac{4^{n-1}n^{n+1}\Gamma\left(n-\frac{1}{2}\right)[\gamma(n,n-1)]^{2}}{\sqrt{\pi}\Gamma(n)^{3}}\right)$ (21)
where, with abuse of notation, we again denoted by a bar this specific term. We also used the result:
$\sum_{k=0}^{n-1}\frac{1}{\Gamma(k+1)^{2}\Gamma(-k+n)^{2}}=\frac{4^{n-1}\Gamma\left(n-\frac{1}{2}\right)}{\sqrt{\pi}\Gamma(n)^{3}}.$
Again apart from the positive numerical term inside the brackets, this term reproduces the correct s-channel factorization for the amplitude with opposite helicities.
Let us quote our proposal for the electromagnetic Compton Scattering amplitude with opposite helicities. Denoting by a bar this particular contribution we find, for a massive particle with spin $s=n>0$ integer:
$\bar{{\cal A}}_{n}({\bf 1},2^{+},{\bf 3},4^{-})=\frac{2ie^{2}}{\left(s-m^{2}\right)\left(u-m^{2}\right)}\mathbb{A}_{n}({\bf 1},2^{+},{\bf 3},4^{-})$ (22)
where
$\mathbb{A}_{n}({\bf 1},2^{+},{\bf 3},4^{-})$$\equiv$ ${\cal N}n^{2}(2n)^{n-1}\gamma(s\alpha^{\prime},n-1)\gamma(u\alpha^{\prime},n-1)$ (23)
$\times$ $\sum_{k=0}^{n-1}\frac{(-1)^{n-1}}{\Gamma(k+1)^{2}\Gamma(-k+n)^{2}}\left(\frac{\bigl{(}q_{2}\cdot\epsilon({\bf 3})\,q_{2}\cdot\epsilon({\bf 1})\bigr{)}}{m^{2}}\right)^{n-1-k}\left(\frac{\bigl{(}q_{4}\cdot\epsilon({\bf 1})\,q_{4}\cdot\epsilon({\bf 3})\bigr{)}}{m^{2}}\right)^{k}$
$\times$ $\epsilon^{+}_{\alpha}(2)\epsilon^{-}_{\beta}(4)\Biggl{\{}-(u-m^{2})\left[\Bigl{(}-\eta^{a_{s}c_{s}}k_{3}^{\beta}+\eta^{a_{s}\beta}q_{4}^{c_{s}}-\eta^{c_{s}\beta}q_{4}^{a_{s}}\Bigr{)}\Bigl{(}k_{3}^{\alpha}+q_{4}^{\alpha}\Bigr{)}\right.$
$+$ $\left.q_{2}^{a_{s}}\left(\eta^{\alpha\beta}q_{4}^{c_{s}}-\eta^{c_{s}\beta}q_{4}^{\alpha}\right)\right.$
$+$ $\left.\eta^{a_{s}\alpha}\Bigl{(}(q_{2}\cdot q_{4})\eta^{c_{s}\beta}-q_{2}^{\beta}q_{4}^{c_{s}}\Bigr{)}+k_{3}^{\beta}\left(q_{2}^{c_{s}}\eta^{a_{s}\alpha}-q_{2}^{a_{s}}\eta^{\alpha c_{s}}\right)\right]$
-- $(s-m^{2})\left[\Bigl{(}-\eta^{a_{s}c_{s}}k_{3}^{\alpha}+q_{2}^{c_{s}}\eta^{a_{s}\alpha}-q_{2}^{a_{s}}\eta^{\alpha c_{s}}\Bigr{)}\Bigl{(}k_{3}^{\beta}+q_{2}^{\beta}\Bigr{)}+q_{4}^{a_{s}}\left(\eta^{\alpha\beta}q_{2}^{c_{s}}-\eta^{\alpha c_{s}}q_{2}^{\beta}\right)\right.$
$+$ $\left.\eta^{a_{s}\beta}\Bigl{(}(q_{2}\cdot q_{4})\eta^{\alpha c_{s}}-q_{2}^{c_{s}}q_{4}^{\alpha}\Bigr{)}+k_{3}^{\alpha}\left(q_{4}^{c_{s}}\eta^{a_{s}\beta}-q_{4}^{a_{s}}\eta^{c_{s}\beta}\right)\right]$
$+$ $\frac{1}{2}\eta^{\alpha\beta}\eta^{a_{s}c_{s}}(s-m^{2})(u-m^{2})\Biggr{\}}\epsilon_{a_{s}}({\bf 1})\epsilon_{c_{s}}({\bf 3}).$
Notice that we have added an arbitrary normalization factor ${\cal N}(n)$ (with ${\cal N}(1)=1$) in order to take care of the multiplicative numerical factor that appears in the factorization of the physical channels.
Let us summarize the three constraints we have imposed:
  • Soft momentum transfer, $q^{2}\lll 1/\alpha^{\prime}$;
  • Compactification of the polarizations $\epsilon^{a_{n}}({\bf 1})$ and $\epsilon^{c_{n}}({\bf 3})$, so that $n\to S$;
  • Consistent factorization in the physical channels for any helicity choice.
[…]
$\epsilon^{\mu}({\bf 3})\epsilon^{\nu}({\bf 1})\to\bar{\epsilon}^{\mu}({\bf 1})\epsilon^{\nu}({\bf 1})=-\langle S^{(\mu}S^{\nu)}\rangle+\frac{i}{2m}\varepsilon^{\mu\nu\rho\sigma}k_{1\rho}\langle S_{\sigma}\rangle-\Theta^{\mu\nu}\equiv\Pi^{\mu\nu}$ (24)
up to subleading terms in the classical limit, where $\Theta^{\mu\nu}$ is the spin-1 projector. For the spin-S case we take appropriate tensor products:
$\bar{\epsilon}^{\mu_{1}}({\bf 1})\cdots\bar{\epsilon}^{\mu_{S}}({\bf 1})\epsilon^{\nu_{1}}({\bf 1})\cdots\epsilon^{\nu_{S}}({\bf 1})=\Pi^{\mu_{1}\nu_{1}}\otimes\Pi^{\mu_{2}\nu_{2}}\otimes\cdots\otimes\Pi^{\mu_{S}\nu_{S}}$ (25)
where a symmetrised expectation value is understood over products of the spin operator. Now one inserts Eq. (25) in Eq. (23) and takes $n\to\infty$. In this limit one may identify $\langle(q\cdot S/m)^{k}\rangle$ with the classical power law $(q\cdot a)^{k}$, where $a^{\mu}$ is the classical ring-radius. We define the resulting expression as being the classical amplitude associated with (23). It is easy to see that, in the classical limit, the summation in (23) will generate powers such as $(q_{2}\cdot a)^{2l}(q_{4}\cdot a)^{2p}$, $l,p$ being integers smaller then $n$, times a piece that is proportional to the spin-1 case. This multiplicative piece satisfies the so-called spin-shift symmetry [22, 25, 43], $a_{i}^{\mu}\to a_{i}^{\mu}+\iota_{i}q^{\mu}/q^{2}$, where $q$ is the momentum transfer and $\iota_{i}$ is an arbitrary parameter. As emphasized by Ref. [25], the emergence of lone factors of $q\cdot a$ will violate the spin-shift symmetry. We conclude that the classical limit of our proposal for the opposite-helicity amplitude should fail to obey this symmetry for $n>1$. However, the massive-spin one ($n=1$ case) opposite-helicity Compton amplitude respects the spin-shift symmetry, as it should.
4 Gravitational Compton amplitudes for all spins
Now let us study the gravity Compton amplitude. Using our previous prescription for deriving the associated field-theory amplitudes, we get
$M^{\alpha\gamma,\beta\delta}_{4}({\bf 1},2,{\bf 3},4)=8i\left(\frac{(-1)^{n}}{n[\Gamma(n)]^{4}}\right)\kappa^{2}m^{2}\frac{1}{\alpha^{\prime}t(s\alpha^{\prime}-4n)(u\alpha^{\prime}-4n)}\left[{\cal A}_{0}^{\alpha\beta}{\cal A}_{0}^{\gamma\delta}\right]_{\alpha^{\prime}\to\alpha^{\prime}/4}$ (26)
where we have expanded $\sin\left(\pi x\right)$ around $x=n$ and also restored the gravity coupling constants, $M_{N}\to\left(\kappa/2\right)^{N-2}M_{N}$ [76], where $\kappa^{2}=32\pi G$. The first mass level $n=1$ corresponds in our prescription to a massive spin-2 particle since here we are compactifying two polarizations of each of the gauge amplitudes. In turn, it is easy to see that our expression agrees with the one derived in Ref. [22] in quadratic order in $a^{\mu}$, as expected from the KLT formula, see also Ref. [77]. In fact, up to quartic order in classical spin our results also match with the ones given in Ref. [22].
After imposing the three constraints listed above, the KLT relations allows us to derive our candidate opposite helicity gravity Compton amplitude which is valid for all quantum spins $S=2n>0$:
¯$\bar{\cal M}_{S}({\bf 1},2^{++},{\bf 3},4^{--})$ $=$ $8i\left(\frac{(-1)^{n}}{n[\Gamma(n)]^{4}}\right)\kappa^{2}m^{2}\frac{1}{\alpha^{\prime}t(\alpha^{\prime}s-4n)(\alpha^{\prime}u-4n)}$ (27)
$\times$ $\left[\alpha^{\prime 4}\mathbb{A}_{S/2}({\bf 1},2^{+},{\bf 3},4^{-})\mathbb{A}_{S/2}({\bf 1},2^{+},{\bf 3},4^{-})\right]_{\alpha^{\prime}\to\alpha^{\prime}/4}.$
One can also similarly construct the all plus gravitational amplitude from the corresponding electromagnetic case and the result matches the one known in the literature [69], apart from a positive multiplicative numerical factor. Moreover, the classical limit of the opposite-helicity amplitude does not satisfy the spin-shift symmetry for quantum spins greater than 2; in any case, the massive-spin 2 opposite-helicity amplitude ($S=2$) obtained from the above double copy formula preserves this symmetry.
We impose the constraints on the gravitational amplitude after performing the double copy procedure on the full string amplitude. This could make the double-copy construction compatible with factorization poles that match the known 3-point amplitudes. We believe this happens because our amplitude inherits some aspects of the full original string amplitude; it is a “stringy” quantum-field-theory amplitude. To confirm this it is also necessary to exploit the R sector of string theory.
5 Conclusions and perspectives
In this work we have proposed candidate Compton amplitudes which are valid to all integer spins. An amusing feature is that the total amplitude (14) contains both Comptons for same and opposite helicities of photons (and gravitons, in the gravitational case, Eq. (26)). This observation is somewhat reminiscent of the Compton amplitude written in terms of massive spinors in the high-energy limit [68]; for instance, for the massive spin-one amplitude, the high-energy limit is given by a sum of its “helicity constituents”. As helicity amplitudes the sum of components is generally not allowed, yet this is precisely how one can sort out the different pieces of the amplitude that unifies all the different helicity amplitudes merged into a single object. Remarkably we observe a similar phenomenon in the present case.
There are other prominent features of such amplitudes which potentially justify the given interpretation. We list these here:
  1. 1.
    Amplitudes are gauge invariant; it is possible to rewrite them in terms of gauge invariant operators;
  2. 2.
    Amplitudes do not present any spurious poles and hence they are valid to all spins;
  3. 3.
    We have found no evidence of the general validity of the spin-shift symmetry;
  4. 4.
    By construction, it has the correct factorization in the physical channels;
  5. 5.
    The gravitational amplitudes obey double copy.
One possible conclusion is that subleading Regge trajectories seem to play an important role here. Indeed, as we have shown in the case of the first excited state $n=1$, the compactification down to the spin-$1$ case allowed us to recover known results in the literature and supplied the motivation to search for a valid amplitude for higher spins based on a string-theory calculation. Perhaps subleading Regge states could be the key to understand why this kind of compactification is needed here.
Our results require further explorations and many of the properties listed above are calling for a better comprehension. For instance, our soft expansion requires some refinement, perhaps following standard procedures in the literature [78, 79, 80]. On the other hand, we need to outline a better comparison with other works on the Compton amplitude such as Refs. [30, 31, 32, 33, 21, 22, 41, 58]. For instance, one could expect that contact terms may look different as in the present case they were fixed by the string excitations. These comparisons are important in order to connect our results with the existing literature and also to understand better the classical limit of our candidate Compton amplitude. Indeed, the worldsheet interpretation for Kerr provided by Ref. [15] might fit well with our description. In turn, as mentioned above, to obtain the corresponding gravitational amplitudes we not only correlate the helicites in the gauge amplitudes, but also the spins of the massive particles. So it is difficult to discuss the independence of the amplitude on the split $S\to S_{1}+S_{2}$ within our formalism. This also motivates the study of the R sector of the superstring. Such studies will be reported in a separate publication elsewhere.
Acknowledgements
We thank Paolo di Vecchia and Francesco Alessio for suggestions and comments on the draft, and for sharing with us their notes with similar calculations and results. We also thank Yilber Fabian Bautista, Radu Roiban, Zvi Bern, N. Emil J. Bjerrum-Bohr, Renann Jusinskas, Rafael Aoude, Riccardo Gonzo, Matteo Sergola, Henrik Johansson and Fei Teng for their invaluable remarks and suggestions. GM acknowledges partial support from FAPESP under grant 2023/06508-8, CNPq under grant 317548/2021-2 and FAPERJ under grant E-26/201.142/2022.
Appendix A. Details of the formula for the Compton scattering amplitude
Here we present the explicit for the quantity ${\cal A}_{0}^{\alpha\beta}$. The amplitude reads
$A^{\alpha\beta}_{4}({\bf 1},2,{\bf 3},4)\equiv A^{\alpha\beta}_{t}=\frac{{\cal V}_{t}}{\alpha^{\prime}t\,\prod_{k=1}^{2n-1}(\alpha^{\prime}t-k)}{\cal A}_{0}^{\alpha\beta}$ (28)
where we have written
${\cal A}_{0}^{\alpha\beta}\equiv{\cal A}^{\alpha\beta(1)}_{4}+{\cal A}^{\alpha\beta(2)}_{4}+{\cal A}^{\alpha\beta(3)}_{4}+{\cal A}^{\alpha\beta(4)}_{4}$ (29)
with
${\cal A}^{\alpha\beta(1)}_{4}({\bf 1},2,{\bf 3},4)$ $=$ $2i\alpha^{\prime}e^{2}(n!)^{2}\sum_{j=0}^{n}\frac{(2\alpha^{\prime})^{n-j}}{j![(n-j)!]^{2}}\eta^{a(j)c(j)}$
$\times$ $\sum_{k=0}^{n-j}\begin{pmatrix}n-j\\ k\end{pmatrix}(-1)^{k}q_{2}^{a(n-j-k)}q_{4}^{a(k)}\sum_{l=0}^{n-j}\begin{pmatrix}n-j\\ l\end{pmatrix}(-1)^{n-j-l}q_{2}^{c(n-j-l)}q_{4}^{c(l)}$
$\times$ $\biggl{\{}\frac{1}{(\alpha^{\prime}t+1)}\gamma(s\alpha^{\prime},k-l+n)\gamma(u\alpha^{\prime},-k+l+n)\eta^{a_{S}c_{S}}\left(q_{2}^{\beta}q_{4}^{\alpha}-(q_{2}\cdot q_{4})\eta^{\alpha\beta}\right)$
$+$ $\Bigl{[}(-u\alpha^{\prime}-k+l+n)\left(q_{2}^{a_{S}}\left(q_{4}^{c_{S}}\eta^{\alpha\beta}-q_{4}^{\alpha}\eta^{c_{S}\beta}\right)-\eta^{a_{S}\alpha}\left(q_{2}^{\beta}q_{4}^{c_{S}}-(q_{2}\cdot q_{4})\eta^{c_{S}\beta}\right)\right)$
-- $(-s\alpha^{\prime}+k-l+n)\left(q_{4}^{a_{S}}\left(q_{2}^{\beta}\eta^{\alpha c_{S}}-q_{2}^{c_{S}}\eta^{\alpha\beta}\right)-\eta^{a_{S}\beta}\left((q_{2}\cdot q_{4})\eta^{\alpha c_{S}}-q_{2}^{c_{S}}q_{4}^{\alpha}\right)\right)\Bigr{]}$
$\times$ $\gamma(s\alpha^{\prime},k-l+n-1)\gamma(u\alpha^{\prime},-k+l+n-1)\biggr{\}}$ (35)
${\cal A}^{\alpha\beta(2)}_{4}({\bf 1},2,{\bf 3},4)$ $=$ $2i\alpha^{\prime}e^{2}(n!)^{2}\sum_{j=0}^{n}\frac{(2\alpha^{\prime})^{n-j}}{j![(n-j)!]^{2}}\eta^{a(j)c(j)}\left(q_{2}^{a_{S}}\eta^{\alpha c_{S}}-q_{2}^{c_{S}}\eta^{a_{S}\alpha}\right)$
$\times$ $\biggl{\{}-\sum_{k=0}^{n-j}\begin{pmatrix}n-j\\ k\end{pmatrix}(-1)^{k}q_{2}^{a(n-j-k)}q_{4}^{a(k)}\sum_{l=0}^{n-j}\begin{pmatrix}n-j\\ l\end{pmatrix}(-1)^{n-j-l}q_{2}^{c(n-j-l)}q_{4}^{c(l)}$
$\times$ $\left((-s\alpha^{\prime}+k-l+n)q_{2}^{\beta}+\alpha^{\prime}t\,k_{3}^{\beta}\right)\gamma(s\alpha^{\prime},k-l+n-1)\gamma(u\alpha^{\prime},-k+l+n-1)$
$+$ $\frac{(n-j)\alpha^{\prime}t}{(2\alpha^{\prime})}\sum_{k=0}^{n-j-1}\begin{pmatrix}n-j-1\\ k\end{pmatrix}(-1)^{k}q_{2}^{a(n-j-1-k)}q_{4}^{a(k)}$ (43)
$\times$ $\sum_{l=0}^{n-j-1}\begin{pmatrix}n-j-1\\ l\end{pmatrix}(-1)^{n-j-1-l}q_{2}^{c(n-j-1-l)}q_{4}^{c(l)}$
$\times$ $\left[\eta^{\beta a_{n}}\Bigl{(}-\gamma(s\alpha^{\prime},k-l+n)\gamma(u\alpha^{\prime},-k+l+n-2)q_{2}^{c_{n}}\right.$
$+$ $\left.\gamma(s\alpha^{\prime},k-l+n-1)\gamma(u\alpha^{\prime},-k+l+n-1)q_{4}^{c_{n}}\Bigr{)}\right.$
$+$ $\left.\eta^{\beta c_{n}}\Bigl{(}\gamma(s\alpha^{\prime},k-l+n-2)\gamma(u\alpha^{\prime},-k+l+n)q_{2}^{a_{n}}\right.$
-- $\left.\gamma(s\alpha^{\prime},k-l+n-1)\gamma(u\alpha^{\prime},-k+l+n-1)q_{4}^{a_{n}}\Bigr{)}\right]\biggr{\}}$ (47)
${\cal A}^{\alpha\beta(3)}_{4}({\bf 1},2,{\bf 3},4)$ $=$ $2i\alpha^{\prime}e^{2}(n!)^{2}\sum_{j=0}^{n}\frac{(2\alpha^{\prime})^{n-j}}{j![(n-j)!]^{2}}\eta^{a(j)c(j)}\left(q_{4}^{c_{S}}\eta^{a_{S}\beta}-q_{4}^{a_{S}}\eta^{c_{S}\beta}\right)$
$\times$ $\left\{\sum_{k=0}^{n-j}\begin{pmatrix}n-j\\ k\end{pmatrix}(-1)^{k}q_{2}^{a(n-j-k)}q_{4}^{a(k)}\sum_{l=0}^{n-j}\begin{pmatrix}n-j\\ l\end{pmatrix}(-1)^{n-j-l}q_{2}^{c(n-j-l)}q_{4}^{c(l)}\right.$
$\times$ $\left.\Bigl{(}(-u\alpha^{\prime}-k+l+n)q_{4}^{\alpha}+\alpha^{\prime}t\,k_{3}^{\alpha}\Bigr{)}\gamma(s\alpha^{\prime},k-l+n-1)\gamma(u\alpha^{\prime},-k+l+n-1)\right.$
$+$ $\left.\frac{(n-j)\alpha^{\prime}t}{(2\alpha^{\prime})}\sum_{k=0}^{n-j-1}\begin{pmatrix}n-j-1\\ k\end{pmatrix}(-1)^{k}q_{2}^{a(n-j-1-k)}q_{4}^{a(k)}\right.$ (55)
$\times$ $\left.\sum_{l=0}^{n-j-1}\begin{pmatrix}n-j-1\\ l\end{pmatrix}(-1)^{n-j-1-l}q_{2}^{c(n-j-1-l)}q_{4}^{c(l)}\right.$
$\times$ $\left.\left[\eta^{\alpha a_{n}}\Bigl{(}-\gamma(s\alpha^{\prime},k-l+n-1)\gamma(u\alpha^{\prime},-k+l+n-1)q_{2}^{c_{n}}\right.\right.$
$+$ $\left.\left.\gamma(s\alpha^{\prime},k-l+n-2)\gamma(u\alpha^{\prime},-k+l+n)q_{4}^{c_{n}}\Bigr{)}\right.\right.$
$+$ $\left.\left.\eta^{\alpha c_{n}}\Bigl{(}\gamma(s\alpha^{\prime},k-l+n-1)\gamma(u\alpha^{\prime},-k+l+n-1)q_{2}^{a_{n}}\right.\right.$
-- $\left.\left.\gamma(s\alpha^{\prime},k-l+n)\gamma(u\alpha^{\prime},-k+l+n-2)q_{4}^{a_{n}}\Bigr{)}\right]\right\}$ (59)
and
${\cal A}^{\alpha\beta(4)}_{4}({\bf 1},2,{\bf 3},4)$ $=$ $2i\alpha^{\prime}e^{2}(n!)^{2}\sum_{j=0}^{n}\frac{(2\alpha^{\prime})^{n-j}}{j![(n-j)!]^{2}}\eta^{a(j)c(j)}$
$\times$ $\left\{-\sum_{k=0}^{n-j}\begin{pmatrix}n-j\\ k\end{pmatrix}(-1)^{k}q_{2}^{a(n-j-k)}q_{4}^{a(k)}\sum_{l=0}^{n-j}\begin{pmatrix}n-j\\ l\end{pmatrix}(-1)^{n-j-l}q_{2}^{c(n-j-l)}q_{4}^{c(l)}\right.$
$\times$ $\left.\biggl{[}\frac{1}{(\alpha^{\prime}t+1)}\gamma(s\alpha^{\prime},k-l+n)\gamma(u\alpha^{\prime},-k+l+n)q_{4}^{\alpha}q_{2}^{\beta}\right.$
$+$ $\left.\Bigl{(}k_{3}^{\alpha}\bigl{(}(-s\alpha^{\prime}+k-l+n)q_{2}^{\beta}+\alpha^{\prime}t\,k_{3}^{\beta}\bigr{)}+(-u\alpha^{\prime}-k+l+n)q_{4}^{\alpha}k_{3}^{\beta}\Bigr{)}\right.$
$\times$ $\left.\gamma(s\alpha^{\prime},k-l+n-1)\gamma(u\alpha^{\prime},-k+l+n-1)\biggr{]}\right.$
-- $\left.\frac{(n-j)}{(2\alpha^{\prime})}\sum_{k=0}^{n-j-1}\begin{pmatrix}n-j-1\\ k\end{pmatrix}(-1)^{k}q_{2}^{a(n-j-1-k)}q_{4}^{a(k)}\right.$ (67)
$\times$ $\left.\sum_{l=0}^{n-j-1}\begin{pmatrix}n-j-1\\ l\end{pmatrix}(-1)^{n-j-1-l}q_{2}^{c(n-j-1-l)}q_{4}^{c(l)}\right.$
$\times$ $\left.\biggl{[}q_{2}^{\beta}\biggl{(}\eta^{\alpha a_{n}}\left(-\gamma(s\alpha^{\prime},k-l+n)\gamma(u\alpha^{\prime},-k+l+n-1)q_{2}^{c_{n}}\right.\right.$
$+$ $\left.\left.\gamma(s\alpha^{\prime},k-l+n-1)\gamma(u\alpha^{\prime},-k+l+n)q_{4}^{c_{n}}\right)\right.$
$+$ $\left.\eta^{\alpha c_{n}}\left(\gamma(s\alpha^{\prime},k-l+n)\gamma(u\alpha^{\prime},-k+l+n-1)q_{2}^{a_{n}}\right.\right.$
-- $\left.\left.\gamma(s\alpha^{\prime},k-l+n+1)\gamma(u\alpha^{\prime},-k+l+n-2)q_{4}^{a_{n}}\right)\biggr{)}\right.$
$+$ $\left.\alpha^{\prime}t\,k_{3}^{\beta}\biggl{(}\eta^{\alpha a_{n}}\left(-\gamma(s\alpha^{\prime},k-l+n-1)\gamma(u\alpha^{\prime},-k+l+n-1)q_{2}^{c_{n}}\right.\right.$
$+$ $\left.\left.\gamma(s\alpha^{\prime},k-l+n-2)\gamma(u\alpha^{\prime},-k+l+n)q_{4}^{c_{n}}\right)\right.$
$+$ $\left.\eta^{\alpha c_{n}}\left(\gamma(s\alpha^{\prime},k-l+n-1)\gamma(u\alpha^{\prime},-k+l+n-1)q_{2}^{a_{n}}\right.\right.$
-- $\left.\left.\gamma(s\alpha^{\prime},k-l+n)\gamma(u\alpha^{\prime},-k+l+n-2)q_{4}^{a_{n}}\right)\biggr{)}\biggr{]}\right.$
$+$ $\left.\frac{(n-j)}{(2\alpha^{\prime})}\sum_{k=0}^{n-j-1}\begin{pmatrix}n-j-1\\ k\end{pmatrix}(-1)^{k}q_{2}^{a(n-j-1-k)}q_{4}^{a(k)}\right.$ (73)
$\times$ $\left.\sum_{l=0}^{n-j-1}\begin{pmatrix}n-j-1\\ l\end{pmatrix}(-1)^{n-j-1-l}q_{2}^{c(n-j-1-l)}q_{4}^{c(l)}\right.$
$\times$ $\left.\biggl{[}\alpha^{\prime}t\,k_{3}^{\alpha}\Bigl{(}\eta^{\beta a_{n}}\left(-\gamma(s\alpha^{\prime},k-l+n)\gamma(u\alpha^{\prime},-k+l+n-2)q_{2}^{c_{n}}\right.\right.$
$+$ $\left.\left.\gamma(s\alpha^{\prime},k-l+n-1)\gamma(u\alpha^{\prime},-k+l+n-1)q_{4}^{c_{n}}\right)\right.$
$+$ $\left.\eta^{\beta c_{n}}\left(\gamma(s\alpha^{\prime},k-l+n-2)\gamma(u\alpha^{\prime},-k+l+n)q_{2}^{a_{n}}\right.\right.$
-- $\left.\left.\gamma(s\alpha^{\prime},k-l+n-1)\gamma(u\alpha^{\prime},-k+l+n-1)q_{4}^{a_{n}}\right)\Bigr{)}\right.$
$+$ $\left.q_{4}^{\alpha}\Bigl{(}\eta^{\beta a_{n}}\left(-\gamma(s\alpha^{\prime},k-l+n)\gamma(u\alpha^{\prime},-k+l+n-1)q_{2}^{c_{n}}\right.\right.$
$+$ $\left.\left.\gamma(s\alpha^{\prime},k-l+n-1)\gamma(u\alpha^{\prime},-k+l+n)q_{4}^{c_{n}}\right)\right.$
$+$ $\left.\eta^{\beta c_{n}}\left(\gamma(s\alpha^{\prime},k-l+n-2)\gamma(u\alpha^{\prime},-k+l+n+1)q_{2}^{a_{n}}\right.\right.$
-- $\left.\left.\gamma(s\alpha^{\prime},k-l+n-1)\gamma(u\alpha^{\prime},-k+l+n)q_{4}^{a_{n}}\right)\Bigr{)}\biggr{]}\right.$
$+$ $\left.\frac{(n-j)(n-j-1)\alpha^{\prime}t}{(2\alpha^{\prime})^{2}}\sum_{k=0}^{n-j-2}\begin{pmatrix}n-j-2\\ k\end{pmatrix}(-1)^{k}q_{2}^{a(n-j-2-k)}q_{4}^{a(k)}\right.$ (79)
$\times$ $\left.\sum_{l=0}^{n-j-2}\begin{pmatrix}n-j-2\\ l\end{pmatrix}(-1)^{n-j-2-l}q_{2}^{c(n-j-2-l)}q_{4}^{c(l)}\right.$
$\times$ $\left.\biggl{[}\eta^{\alpha a_{n}}\eta^{\beta a_{n-1}}\Bigl{(}\gamma(s\alpha^{\prime},k-l+n)\gamma(u\alpha^{\prime},-k+l+n-2)q_{2}^{c_{n-1}}q_{2}^{c_{n}}\right.$
-- $\left.\gamma(s\alpha^{\prime},k-l+n-1)\gamma(u\alpha^{\prime},-k+l+n-1)q_{2}^{c_{n-1}}q_{4}^{c_{n}}\right.$
-- $\left.\gamma(s\alpha^{\prime},k-l+n-1)\gamma(u\alpha^{\prime},-k+l+n-1)q_{4}^{c_{n-1}}q_{2}^{c_{n}}\right.$
$+$ $\left.\gamma(s\alpha^{\prime},k-l+n-2)\gamma(u\alpha^{\prime},-k+l+n)q_{4}^{c_{n-1}}q_{4}^{c_{n}}\Bigr{)}\right.$
$+$ $\left.\eta^{\alpha c_{n}}\eta^{\beta c_{n-1}}\Bigl{(}\gamma(s\alpha^{\prime},k-l+n-2)\gamma(u\alpha^{\prime},-k+l+n)q_{2}^{a_{n-1}}q_{2}^{a_{n}}\right.$
-- $\left.\gamma(s\alpha^{\prime},k-l+n-1)\gamma(u\alpha^{\prime},-k+l+n-1)q_{2}^{a_{n-1}}q_{4}^{a_{n}}\right.$
-- $\left.\gamma(s\alpha^{\prime},k-l+n-1)\gamma(u\alpha^{\prime},-k+l+n-1)q_{4}^{a_{n-1}}q_{2}^{a_{n}}\right.$
$+$ $\left.\gamma(s\alpha^{\prime},k-l+n)\gamma(u\alpha^{\prime},-k+l+n-2)q_{4}^{a_{n-1}}q_{4}^{a_{n}}\Bigr{)}\biggr{]}\right.$
$+$ $\left.\frac{(n-j)^{2}\alpha^{\prime}t}{(2\alpha^{\prime})^{2}}\sum_{k=0}^{n-j-1}\begin{pmatrix}n-j-1\\ k\end{pmatrix}(-1)^{k}q_{2}^{a(n-j-1-k)}q_{4}^{a(k)}\right.$ (85)
$\times$ $\left.\sum_{l=0}^{n-j-1}\begin{pmatrix}n-j-1\\ l\end{pmatrix}(-1)^{n-j-1-l}q_{2}^{c(n-j-1-l)}q_{4}^{c(l)}\right.$
$\times$ $\left.\Bigl{[}\gamma(s\alpha^{\prime},k-l+n-2)\gamma(u\alpha^{\prime},-k+l+n)\eta^{\alpha a_{n}}\eta^{\beta c_{n}}\right.$
$+$ $\left.\gamma(s\alpha^{\prime},k-l+n)\gamma(u\alpha^{\prime},-k+l+n-2)\eta^{\alpha c_{n}}\eta^{\beta a_{n}}\Bigr{]}\right.$
$+$ $\left.\frac{1}{(2\alpha^{\prime})}\eta^{\alpha\beta}\sum_{k=0}^{n-j}\begin{pmatrix}n-j\\ k\end{pmatrix}(-1)^{k}q_{2}^{a(n-j-k)}q_{4}^{a(k)}\sum_{l=0}^{n-j}\begin{pmatrix}n-j\\ l\end{pmatrix}(-1)^{n-j-l}q_{2}^{c(n-j-l)}q_{4}^{c(l)}\right.$
$\times$ $\left.(-s\alpha^{\prime}+k-l+n)(-u\alpha^{\prime}-k+l+n)\right.$
$\times$ $\left.\frac{1}{(\alpha^{\prime}t+1)}\gamma(s\alpha^{\prime},k-l+n-1)\gamma(u\alpha^{\prime},-k+l+n-1)\right\}\eta^{a_{S}c_{S}}$ (94)
where $\gamma(x,N)\equiv\prod_{m=1}^{N}(-x+m)=(1-x)_{N}$, $q^{a(m)}\equiv q^{a_{1}}q^{a_{2}}\cdots q^{a_{m}}$, etc., and also we defined $\eta^{a(j)c(j)}\equiv\eta^{a_{1}c_{1}}\eta^{a_{2}c_{2}}\cdots\eta^{a_{j}c_{j}}$. It may appear that there is a tachyonic pole, but if we group all such contributions together:
$\frac{1}{(\alpha^{\prime}t+1)}\left[\eta^{a_{S}c_{S}}\left(q_{2}^{\beta}q_{4}^{\alpha}-(q_{2}\cdot q_{4})\eta^{\alpha\beta}\right)-\eta^{a_{S}c_{S}}\left(q_{2}^{\beta}q_{4}^{\alpha}-\frac{1}{(2\alpha^{\prime})}\eta^{\alpha\beta}\right)\right]$
$=\frac{1}{(\alpha^{\prime}t+1)}\left[\eta^{a_{S}c_{S}}\left(q_{2}^{\beta}q_{4}^{\alpha}+\frac{1}{2\alpha^{\prime}}\alpha^{\prime}t\eta^{\alpha\beta}\right)-\eta^{a_{S}c_{S}}\left(q_{2}^{\beta}q_{4}^{\alpha}-\frac{1}{(2\alpha^{\prime})}\eta^{\alpha\beta}\right)\right]$
$=\frac{1}{(2\alpha^{\prime})(\alpha^{\prime}t+1)}(\alpha^{\prime}t+1)\eta^{\alpha\beta}\eta^{a_{S}c_{S}}=\frac{1}{(2\alpha^{\prime})}\eta^{\alpha\beta}\eta^{a_{S}c_{S}}$ (95)
and so there are no tachyons in the spectrum, as expected.
The other color-ordered amplitudes are given by
$A^{\alpha\beta}_{s}=\frac{{\cal V}_{s}}{\alpha^{\prime}t\,\prod_{k=1}^{2n-1}(\alpha^{\prime}s-k)}{\cal A}_{0}^{\alpha\beta}$ (96)
and
$A^{\alpha\beta}_{u}=\frac{{\cal V}_{u}}{\alpha^{\prime}t\,\prod_{k=1}^{2n-1}(\alpha^{\prime}u-k)}{\cal A}_{0}^{\alpha\beta}$ (97)
where
$A_{t}$$\equiv$ $A_{4}({\bf 1},2,{\bf 3},4)$
$A_{s}$$\equiv$ $A_{4}(2,{\bf 3},{\bf 1},4)$
$A_{u}$$\equiv$ $A_{4}({\bf 3},{\bf 1},2,4).$ (98)
To obtain such expressions we made use of the definitions of ${\cal V}_{t},\gamma(x,N)$ and repeated application of the recurrence formula
$\Gamma(z)=\frac{\Gamma(z+n+1)}{z(z+1)\cdots(z+n)}.$
Therefore, the full non-Abelian amplitude can be written as [71]
$A_{4}^{\textrm{full}}({\bf 1},2,{\bf 3},4)$ $=$ $\textrm{tr}\bigl{[}T^{a_{1}}T^{a_{2}}T^{a_{3}}T^{a_{4}}+(-1)^{2n}T^{a_{4}}T^{a_{3}}T^{a_{2}}T^{a_{1}}\bigr{]}A_{t}$ (99)
$+$ $\textrm{tr}\bigl{[}T^{a_{2}}T^{a_{3}}T^{a_{1}}T^{a_{4}}+(-1)^{2n}T^{a_{4}}T^{a_{1}}T^{a_{3}}T^{a_{2}}\bigr{]}A_{s}$
$+$ $\textrm{tr}\bigl{[}T^{a_{3}}T^{a_{1}}T^{a_{2}}T^{a_{4}}+(-1)^{2n}T^{a_{4}}T^{a_{2}}T^{a_{1}}T^{a_{3}}\bigr{]}A_{u}$
where $T^{a}$ are Chan-Paton matrices, the factor $(-1)^{2n}$ arises due to worldsheet parity. One can use worldsheet monodromy relations,
$\sin(\alpha^{\prime}\pi t)A_{s}$ $=$ $\sin(\alpha^{\prime}\pi s)A_{t}$
$\sin(\alpha^{\prime}\pi u)A_{t}$ $=$ $\sin(\alpha^{\prime}\pi t)A_{u}$
$\sin(\alpha^{\prime}\pi s)A_{u}$ $=$ $\sin(\alpha^{\prime}\pi u)A_{s},$ (100)
in order to obtain relations between the three color-ordered partial amplitudes given above. Finally, the momentum kernel ${\cal S}_{\alpha^{\prime}}$ appearing in the KLT formula given in the main text is given by [75]
${\cal S}_{\alpha^{\prime}}[i_{1},\ldots,i_{k}|j_{1},\ldots j_{k}]_{p}\equiv\left(\frac{\pi\alpha^{\prime}}{2}\right)^{-k}\prod_{t=1}^{k}\sin\left[\frac{\pi\alpha^{\prime}}{2}\left(p\cdot k_{i_{t}}+\sum_{q>t}^{k}\theta(i_{t},i_{q})k_{i_{t}}\cdot k_{i_{q}}\right)\right]$ (101)
where $\theta(i_{t},i_{q})$ equals 1 if the ordering of the legs $i_{t}$ and $i_{q}$ is opposite in the sets $\{i_{1},\ldots,i_{k}\}$ and $\{j_{1},\ldots,j_{k}\}$, and zero if the ordering is the same.

### v2

@@
 1 Introduction
-An important insight arising from studies of quantum scattering amplitudes is that gravitational amplitudes are simpler than one would expect [1].
-The double copy prescription determines that the gravitational amplitudes are obtained as a product of two Yang-Mills amplitudes [2, 3, 4].
+An important insight arising from contemporary studies of quantum scattering amplitudes is that gravitational amplitudes are simpler than one would expect from a brute-force Feynman diagram calculation [1].
+The double copy prescription determines that the gravitational amplitudes are obtained as a product of two Yang-Mills amplitudes.
 The original formulation by Kawai, Lewellen and Tye states that a closed string amplitude at the tree level is given by a sum over the products of two open string amplitudes at tree level [2].
-The current trend is invaluable for the physics of gravitational waves [6, 7, 8, 9] due to the birth of gravitational wave astronomy, driven by landmark observations made by the LIGO and Virgo collaborations [10, 11, 12, 13, 14].
+More recently, Bern, Carrasco and Johansson demonstrated that gravitational numerators are, in a sense, the square of the kinematic numerators of the Yang-Mills theory [3, 4, 5, 6], which has revealed another aspect of the double copy.
+This approach also connects classical solutions of Yang-Mills theory and gravity.
+In particular, point charges in Yang-Mills theory map to point sources in gravity [7, 8, 9, 10].
+There are also indications that the double copy may encompass bound states and particles with spin [11, 12], and should also be relevant for higher-derivative gravity theories [13, 14, 15, 16, 17, 18, 19].
+This current trend can be expected to be of extreme importance for the physics of gravitational waves.
+Indeed, the broader interest in this program is to scrutinize the application of modern scattering amplitude techniques to address the two-body gravity problem.
+This question arises naturally due to the birth of gravitational wave astronomy, driven by landmark observations made by the LIGO and Virgo collaborations [20, 21, 22, 23, 24].
+Such techniques complement several methods presented in the literature and have been of utmost significance in fostering interdisciplinary collaborations, including the exploration of new research directions.
+Amplitudes have already been successfully applied to understand aspects of the general relativistic two-body problem [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 6, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78].
+In fact, in recent years there has been a growing interest in amplitude calculations applied to classical gravity.
+And the benefit is not only on the technical side, but also in obtaining a deeper knowledge that field theory provides for long-standing problems such as interpretation of infrared divergences [79] and the connection between effective potential and scattering angle [80, 81, 82].
+The use of effective field theory methods to classical gravity were investigated in a pioneering way by [83], see also [84, 85, 86, 87, 88, 89, 79, 90, 91].
+They have been successfully applied to the general problem of two-body relativistic dynamics in the post-Newtonian (PN) approximation up to the fourth perturbative order, with partial results in the fifth perturbative order [92, 93].
+The fourth Post-Minkowskian (PM) Order was studied in Ref.
+[94], and, within amplitude-based methods, in Refs.
+[41, 42].
+In particular, in [95] the connection between the classical two-body gravitational problem and the calculation of quantum field theory amplitudes was highlighted, showing how the evaluation of the effective potential of $n$-th order in PN approximation can be interpreted in terms of a calculation of quantum self-energy in diagrams of massless particles with multiple self-interactions.
+Solving the equations iteratively in perturbative orders, as usually done in PN approximation, is reminiscent of the application of perturbative methods in quantum field theory, where scattering amplitude calculations are traditionally organized in Feynman diagrams.
 In this context, we have witnessed a particularly intense flurry of activities about the gravitational interaction of spinning particles [96, 97, 98, 62, 55, 99, 72, 100, 74, 101, 102, 69, 103, 104, 105, 106, 107, 108, 109, 110, 73, 111, 112, 113, 114, 115, 116, 71, 75, 78, 76, 77, 117, 118, 119, 120, 70, 121, 122, 123, 124, 125, 126, 127, 128, 129]  [^4] For a much larger body of older and current research please see the references within [97, 98, 62, 55, 99, 72, 100, 74, 101, 102, 69, 103, 104, 105, 106, 107, 108, 109, 110, 73, 111, 112, 113, 114, 115, 116, 71, 75, 78, 76, 77, 117, 118, 119, 120, 70, 121, 122, 123, 124, 125, 126, 127, 128, 129]..
-Indeed, gravitational dynamics of spinning objects has been a subject of vigorous study for several years [61, 62, 63, 64].
-An outstanding problem emerging in this regard is the evaluation of the gravitational Compton scattering amplitude – the opposite-helicity version is a key ingredient in the evaluation of the conservative potential of compact binary systems with spin [41, 42, 43, 65, 66] and also waveform calculations with spin [60, 59, 67].
-However, previous calculations in the literature displayed the presence of non-physical poles for massive particles with quantum spins greater than one, in the electromagnetic case, and greater than two, in the gravitational case [68, 69]  [^5] More precisely, the opposite-helicity Compton amplitude built from BCFW recursion relations have unphysical poles because this procedure is no longer predictive in the case of higher spins.
-This is due to the fact that the product of BCFW shifted three-point amplitudes does not approach zero fast enough as the shift parameter goes to infinity [25].
+Indeed, gravitational dynamics of spinning objects has been a subject of vigorous study for several years [130, 131, 132, 133] in both PN and PM approaches.
+Since scattering amplitudes are envisaged as useful means to incorporate black-hole dynamics, the expectation is that gravitational dynamics of Kerr black holes can be encoded in the dynamics described by a set of higher-dimension operators with presumably known (or to be inferred) Wilson coefficients.
+Early work suggested that the amplitudes involving massive low-spin states and gravitons could be used to calculate the lowest orders in multipolar spin expansion from black-hole observables, such as scattering angles and gravitational potentials.
+In fact, conservative observables for Kerr black holes have been studied in a variety of different frameworks, including higher-spin effective field theories and worldline field theories [71, 120, 134, 135, 136].
+An outstanding problem emerging in this regard is the evaluation of the gravitational Compton scattering amplitude – the opposite-helicity version is a key ingredient in the evaluation of the conservative potential of compact binary systems with spin [116, 71, 75, 137, 138] and also waveform calculations with spin [129, 128, 139, 140].
+Previous calculations in the literature have shown how challenging this task can be [141, 142]  [^5] Opposite-helicity Compton amplitude built from BCFW recursion relations have unphysical poles because this procedure is no longer predictive in the case of higher spins – the product of BCFW shifted three-point amplitudes does not approach zero fast enough as the shift parameter goes to infinity [102].
 Nevertheless, we remark that if one starts with a local Lagrangian the amplitudes calculated from it using standard Feynman rules are expected to be manifestly local; see Ref.
 [116] for a calculation for the case of massive spin 3 electromagnetic Compton..
-Important ongoing research is being conducted to address this problem [28, 25, 21, 22, 27, 29, 30, 31, 32, 33, 58].
-This scenario opens up a great opportunity to investigate string amplitudes.
-It is known that one of the most notable features of string theory is the presence of an infinite tower of massive spinning modes in its spectrum required for consistency [70, 71].
-However, some difficulties have been highlighted in the literature [72, 73, 42]; in particular, the classical limit of 3-point string amplitudes associated with leading Regge states of the open and closed superstring differ from the expected result for a Kerr black hole.
-This seems to indicate that using string theory to describe Kerr black holes is at the very least a formidable and challenging task.
-In this paper we wish to report significant progress.
-As an unexpected result we extract from a string-theory based calculation the correct tree-level Compton amplitude for a Kerr black hole for quantum spin two – the analogous result for $\sqrt{\textrm{Kerr}}$ is for quantum spin one, which we also recover here  [^6] Paolo di Vecchia and Francesco Alessio have also obtained this result through a slightly different calculation.
-We thank them for sharing their notes with us..
-Having succeeded in recovering this important benchmark, we propose a candidate Compton amplitude which is free from any spurious poles and valid to any integer quantum spin.
-In order to obtain such amplitudes we impose three plausible physical constraints, to be discussed later on in this paper.
-With these simple rules we manage to obtain sensible field-theory Compton amplitudes which are physically sound.
-We are employing natural units, with $\hbar=c=1$.
+In any case, subsequent research efforts have successfully navigated initial obstructions and by now important ongoing research is being conducted to address this problem [104, 102, 72, 100, 103, 105, 106, 107, 108, 109, 127, 143, 144].
+This scenario also opens up a great opportunity to investigate string amplitudes.
+Indeed, one of the most notable features of string theory is the presence of an infinite tower of massive spinning modes in its spectrum, required for consistency.
+Moreover, the string theory framework allows one to consider the scattering of states of indefinite spin in a systematic manner [145].
+However, some difficulties have been highlighted in the literature [146, 147, 71].
+Nevertheless, as string theory provides a consistent theory of massive interacting higher-spin particles, one can regard it as a natural framework to look into classical objects which involve spin-to-infinity limits.
+Indeed, in this paper we wish to report progress in that direction.
+First, as a promissing result, we extract from a string-theory based calculation the correct tree-level Compton amplitude for a Kerr black hole for quantum spin two  [^6] Paolo di Vecchia and Francesco Alessio have also obtained the same result through a slightly different calculation.
+We thank them for sharing their unpublished notes with us..
+Then, having succeeded in recovering this result, we propose a candidate Compton amplitude which is free from any spurious poles and valid to any integer quantum spin.
+In summary, our claim is: string-theory amplitudes could potentially contain amplitudes that describe compact astrophysical bodies, in the form of specific contributions (among many others) in disguise inside the full amplitude.
+To extract those hidden contributions, all other contributions must be projected out, and we propose that this can be achieved by imposing the following three constraints, to be explained in detail in this paper:
+  • soft momentum transfer;
+  • compactification of polarizations, so that the mass level matches the spin of the leading Regge states;
+  • consistent factorization in the physical channels for any helicity choice.
+Perhaps our procedure is best justified a posteriori, given the results we obtain.
+With the simple rules above we manage to obtain sensible field-theory Compton amplitudes which are physically sound and whose classical limit agrees with that of Ref.
+[106] up to fifth order in the ring radius (in the gravitational case).
+In fact, we argue that such a reasoning can also be used to extract the correct 3-point root-Kerr amplitude from the corresponding superstring amplitude [145, 146], up to terms of order (mass)-1.
+Throughout the paper, we employ natural units, with $\hbar=c=1$.
 2 Compton scattering for open and closed strings
@@
 Let us calculate the Compton amplitude involving two photons and two massive spin-$S$ particles.
-For simplicity in this paper we will only study the NS sector of the string.
+For simplicity in this paper we will only study leading Regge states of the NS sector of the superstring.
 We consider the Minkowski metric as $\eta_{\mu\nu}=\textrm{diag}(-,+,+,\cdots,+)$ and Mandelstam variables defined as $s_{ij}=-(k_{i}+k_{j})^{2}$, with $s_{12}=s$, $s_{13}=t$ and $s_{14}=u$.
@@
 $V_{0}^{(0)}(q;z)=e\sqrt{2\alpha^{\prime}}\epsilon_{\nu}(q):\left((q\cdot\psi)\psi^{\nu}(z)+\frac{1}{2\alpha^{\prime}}i\partial X^{\nu}\right)e^{iq\cdot X(z)}:.$ (2)
+See also Ref.
+[148] for the vertex operators of the open bosonic string.
 We employ a given representation of the massive polarization tensors $\epsilon_{(S)}$ (they are symmetric, traceless and transverse in all indices) of spin-S particles in which they are written as symmetric tensor products of massive spin-$1$ polarization vectors $\epsilon_{1}$, namely $\epsilon_{(S)}=\bigotimes_{S}\epsilon_{1}$.
@@
 […]
-where the factor $e^{-\lambda}$ comes from the Euler-number term in the action. For convenience we write the amplitude as
+where the factor $e^{-\lambda}$ comes from the Euler-number term in the action. The above bold notation refers to massive spinning particles. For convenience we write the amplitude as
 $A_{4}({\bf 1},2^{h_{2}},{\bf 3},4^{h_{4}})$ $=$ $\epsilon_{\alpha}(2)\epsilon_{\beta}(4)\,\epsilon({\bf 3})\cdot A^{\alpha\beta}_{4}({\bf 1},2^{h_{2}},{\bf 3},4^{h_{4}})\cdot\epsilon({\bf 1})$ (4)
@@
 Henceforth we are suppressing the higher-spin indices for clarity.
-Moreover, we consider $z_{1}>z_{2}>z_{4}$.
-This means that we integrate $z_{3}$ over $[z_{4},z_{2}]$ to ensure cyclic ordering $(1,2,3,4)$ of the insertion points.
-Here we take $z_{1}\to\infty$, $z_{2}=1$ and $z_{4}=0$, which implies that $z_{3}\in[0,1]$.
+Moreover, we take $z_{1}\to\infty$, $z_{2}=1$ and $z_{4}=0$, which implies that $z_{3}\in[0,1]$.
 We proceed to calculate the integrals by employing well known correlators in the literature [145, 149].
@@
 For the Compton amplitude, the KLT relation reads
-$M_{4}({\bf 1}^{S},2^{h_{2}\tilde{h}_{2}},{\bf 3}^{S},4^{h_{4}\tilde{h}_{4}})$ $=$ $(-1)^{n}\left(\frac{\pi\alpha^{\prime}}{2}\right)^{-1}\sin\left(\frac{\pi\alpha^{\prime}s}{4}\right)$ (8)
+$M_{4}({\bf 1}^{S},2^{h_{2}\tilde{h}_{2}},{\bf 3}^{S},4^{h_{4}\tilde{h}_{4}})$ $=$ $(-1)^{n}\left(\frac{\pi\alpha^{\prime}}{4}\right)^{-1}\sin\left(\frac{\pi\alpha^{\prime}s}{4}\right)$ (8)
 $\times$ $A_{4}({\bf 1}^{S/2},2^{h_{2}},{\bf 3}^{S/2},4^{h_{4}})\bigg{|}_{\alpha^{\prime}\to\alpha^{\prime}/4}\tilde{A}_{4}({\bf 1}^{S/2},2^{\tilde{h}_{2}},4^{\tilde{h}_{4}},{\bf 3}^{S/2})\bigg{|}_{\alpha^{\prime}\to\alpha^{\prime}/4}$
@@
 In our previous notation, we can write the amplitude as
-$M^{\alpha\gamma,\beta\delta}_{4}({\bf 1},2,{\bf 3},4)$ $=$ $(-1)^{n}\left(\frac{\pi\alpha^{\prime}}{2}\right)^{-1}\sin\left(\frac{\pi\alpha^{\prime}s}{4}\right)\sin\left(\frac{\pi\alpha^{\prime}u}{4}\right)$ (9)
+$M^{\alpha\gamma,\beta\delta}_{4}({\bf 1},2,{\bf 3},4)$ $=$ $(-1)^{n}\left(\frac{\pi\alpha^{\prime}}{4}\right)^{-1}\sin\left(\frac{\pi\alpha^{\prime}s}{4}\right)\sin\left(\frac{\pi\alpha^{\prime}u}{4}\right)$ (9)
 $\times$ $\left\{\frac{{\cal V}_{t}}{\sin(\alpha^{\prime}\pi t)\alpha^{\prime}t\,\prod_{k=1}^{2n-1}(\alpha^{\prime}t-k)}\frac{{\cal V}_{t}}{\alpha^{\prime}t\,\prod_{l=1}^{2n-1}(\alpha^{\prime}t-l)}{\cal A}_{0}^{\alpha\beta}{\cal A}_{0}^{\gamma\delta}\right\}_{\alpha^{\prime}\to\alpha^{\prime}/4}$
@@
 3 Electromagnetic Compton amplitudes for all spins
+3.1 Soft expansion
 Our aim is to extract from the above string results terms potentially relevant for the classical description of spinning compact binary systems.
@@
 [146] and define suitable four-dimensional polarizations and momenta.
-Moreover, we also isolate the poles corresponding to the mass level $n$; these are connected to the s-channel and the u-channel.
-To extract the associated field-theory amplitudes for generic spins, we consider both momenta associated with the external photons as being soft, $q\to\hbar q$.
-Observe that keeping only the level $n$ poles corresponds to the leading 1/$1/\hbar$ term in the soft expansion.
-Then at leading order in a soft expansion we find
-$A^{\alpha\beta}_{t}\Bigg{|}_{\textrm{field theory}}\equiv{\cal A}^{\alpha\beta}_{S}=\frac{1}{[\Gamma(n)]^{2}}\frac{{\cal A}_{0}^{\alpha\beta}}{(s\alpha^{\prime}-n)(u\alpha^{\prime}-n)}+{\cal O}(\alpha^{\prime}t).$ (10)
+Moreover, we also isolate the poles corresponding to the mass level $n$; these are connected to the s-channel and the u-channel (since we consider electromagnetic amplitudes, the t-channel is absent).
+To extract our candidate field-theory amplitudes for generic spins, we begin by considering both momenta associated with the external photons as being soft, $q\to\hbar q$.
+To follow this procedure, we must conveniently rewrite the amplitude by using that
+$\frac{\Gamma(1-s\alpha^{\prime})\Gamma(1-u\alpha^{\prime})}{\Gamma(1+\alpha^{\prime}t-2n)}\frac{1}{\alpha^{\prime}t\,\prod_{k=1}^{2n-1}(\alpha^{\prime}t-k)}$ $=$ $\frac{\Gamma(1+2\alpha^{\prime}\,k_{1}\cdot q_{2})}{(1-s\alpha^{\prime})(2-s\alpha^{\prime})\cdots(n-s\alpha^{\prime})}$ (10)
+$\times$ $\frac{\Gamma(1+2\alpha^{\prime}\,k_{1}\cdot q_{4})}{(1-u\alpha^{\prime})(2-u\alpha^{\prime})\cdots(n-u\alpha^{\prime})}$
+$\times$ $\frac{1}{\Gamma(1+\alpha^{\prime}t)}$
+where the recurrence formula
+$\Gamma(z)=\frac{\Gamma(z+n+1)}{z(z+1)\cdots(z+n)}$
+as well as the result
+$\Gamma(z-n)=(-1)^{n-1}\frac{\Gamma(-z)\Gamma(1+z)}{\Gamma(n+1-z)}$
+were used. Observe that keeping only the level $n$ poles corresponds to the leading 1/$1/\hbar$ term in the soft expansion. Then at leading order in a soft expansion we find
+$A^{\alpha\beta}_{t}\Bigg{|}_{\textrm{field theory}}\equiv{\cal A}^{\alpha\beta}_{S}=\frac{1}{[\Gamma(n)]^{2}}\frac{{\cal A}_{0}^{\alpha\beta}}{(s\alpha^{\prime}-n)(u\alpha^{\prime}-n)}.$ (11)
 where the subscript $S$ in ${\cal A}^{\alpha\beta}_{S}$ denotes the spin of the massive particle.
+3.2 First mass level
 Before continuing the discussion for generic spins, let us study the lowest-spin massive boson.
@@
 -- $\frac{q_{2}\cdot q_{4}}{8}\bigl{(}f_{2\alpha\beta}f_{4\lambda\kappa}+f_{4\alpha\beta}f_{2\lambda\kappa}\bigr{)}\{{\cal J}^{\alpha\beta},{\cal J}^{\lambda\kappa}\}\biggr{\}}$
-where $h_{2},h_{4}$ are the helicities of the photons and $F_{i}^{\mu\nu}=2\epsilon^{[\mu}(i)q_{i}^{\nu]}$.
+where $h_{2},h_{4}$ are the helicities of the photons and $f_{i}^{\mu\nu}=2q_{i}^{[\mu}\epsilon^{\nu]}(i)$.
 To obtain this expression we wrote $\epsilon_{a_{S}}({\bf 1})\epsilon_{c_{S}}({\bf 3})={\cal J}_{a_{S}c_{S}}/2+\epsilon_{(a_{S}}({\bf 1})\epsilon_{c_{S})}({\bf 3})$, where we have defined ${\cal J}^{\alpha\gamma}\equiv-i\epsilon^{\mu}({\bf 1})(M^{\alpha\gamma})_{\mu\nu}\epsilon^{\nu}({\bf 3})$, $(M^{ab})_{c}^{\ d}=2i\delta^{[a}_{\ c}\eta^{b]d}$ being the Lorentz generator in the spin-$1$ representation.
@@
 Up to a numerical overall factor, the result (12) perfectly matches the one previously derived by others [100] (except for possible sign differences due to different conventions on momenta).
-By rewriting this expression with the help of spinor-helicity methods, we also find that this amplitude agrees with other known expressions in the literature [68, 69].
-Observe that $\alpha^{\prime}$ is effectively parameterizing the expansion in the soft photon momenta.
-Nevertheless, to see that the approach of taking directly $\alpha^{\prime}\to 0$ is not, in general, the proper way to obtain the field theory limit which is useful in the present context, let us quote the amplitude for the spin-2 leading Regge state.
+The spin multipole decomposition of the Compton amplitude given by (12) can be easily rewritten in the massive spinor-helicity basis [151].
+For the opposite helicity amplitude, $h_{2}=+,h_{4}=-$, one finds that
+$A_{S}({\bf 1},2^{+},{\bf 3},4^{-})$ $=$ $-\frac{2ie^{2}}{m^{2S}}\frac{\langle 4|{\bf 1}|2\bigr{]}^{2}}{(s-m^{2})(u-m^{2})}\langle{\bf 3}|^{2S}\exp\left(i\frac{q_{4}^{\mu}\bar{\epsilon}_{4}^{\nu}{\cal J}_{1\mu\nu}}{\bar{p}\cdot\bar{\epsilon}_{4}}\right)|{\bf 1}\rangle^{2S}$ (14)
+$=$ $-\frac{2ie^{2}}{(s-m^{2})(u-m^{2})}\langle 4|{\bf 1}|2\bigl{]}^{2-2S}\Bigl{(}\langle{\bf 3}4\rangle\bigl{[}{\bf 1}2\bigr{]}+\langle{\bf 1}4\rangle\bigl{[}{\bf 3}2\bigr{]}\Bigr{)}^{2S}\,\,\,\,\,\,\,\,(S\leq 1)$
+where $\bar{p}=(p_{1}-p_{3})/2$ and $\bar{\epsilon}_{4}$ is the polarization vector for $q_{4}$ with reference momentum fixed to be $q_{2}$, see also [113].
+A similar calculation can be performed for the same-helicity amplitude.
+Therefore the spin-1 amplitude, computed from a string-theory based calculation, agrees with other known expressions in the literature [141, 142].
+Observe from Eq.
+(10) that $\alpha^{\prime}$ is effectively parameterizing the expansion in the soft photon momenta, since it always appears accompanied by a Mandelstam variable.
+Nevertheless, to see that the approach of taking directly $\alpha^{\prime}\to 0$ is not, in general, the proper way to obtain the field theory limit which is useful in the present context, let us quote the amplitude for the spin-2 leading Regge state with this prescription.
 The calculation is rather involved but straightforward.
@@
 Besides, strictly speaking, the $\alpha^{\prime}\to 0$ amounts to projecting out all massive spinning particles from the spectrum.
-These are clearly indications that the limit $\alpha^{\prime}\to 0$ is not adequate to our purposes.
-Given the precedent successful computation, Eq.
-(11), let us proceed to the general case.
+These are clearly indications that the straightforward limit $\alpha^{\prime}\to 0$ is not adequate to our purposes.
+3.3 All mass levels
+Given the precedent successful computation of the amplitude for the massive spin-$1$ case, Eq.
+(12), we now undertake the task of investigating the general case, with the polarization tensors given by $\epsilon_{a_{1}\cdots a_{S}}({\bf 1})=\epsilon_{a_{1}}({\bf 1})\ldots\epsilon_{a_{n}}({\bf 1})\epsilon_{a_{S}}({\bf 1})$ and $\epsilon_{c_{1}\cdots c_{S}}({\bf 3})=\epsilon_{c_{1}}({\bf 3})\ldots\epsilon_{c_{n}}({\bf 3})\epsilon_{c_{S}}({\bf 3})$.
 Following previous prescriptions, we consider the compactification of two massive polarization vectors, namely $\epsilon_{a_{n}}({\bf 1})$ and $\epsilon_{c_{n}}({\bf 3})$.
@@
 The number of the mass level $n$ characterizes the spin of the massive particle under consideration and we can just take $n\to S$ in the above expressions.
-There is still one more constraint that we should consider – we require the associated residues to be given by the product of two minimally coupled 3-point amplitudes.
+Moreover, we remark that this prescription is valid to all mass levels $n\geq 1$.
+The Compton amplitude given by Eq.
+(16) is the full quantum-field-theory amplitude after imposing the two aforementioned constraints, that is, messengers with soft momenta and adequate compactification of one of the massive polarization vectors.
+As we can see, there are many contributions to the full amplitude, many of them are unlikely to play a role in the root-Kerr problem.
+However, we will now make the following conjecture: The amplitude whose classical limit describes the dynamics of a compact body is secretly hidden in the full quantum amplitude (16) as a specific contribution.
+To expose such a contribution, it is obvious that the two requirements established above are not enough; instead, there should be at least one extra step.
+We now require the associated residues to be given by the product of two minimally coupled 3-point amplitudes.
 For this purpose, let us analyze the s-channel cut (similar considerations apply to the u-channel cut).
@@
 First let us consider the all plus amplitude.
-The term in square brackets is just the spin multipole decomposition of the QED Compton amplitude for matter of spin one [22], so this must be proportional to $\bigl{[}24\bigr{]}^{2}\langle{\bf 1}{\bf 3}\rangle^{2}/t$ [69].
+The term in square brackets is just the spin multipole decomposition of the QED Compton amplitude for matter of spin one [100], so this must be proportional to $\bigl{[}24\bigr{]}^{2}\langle{\bf 1}{\bf 3}\rangle^{2}/t$ in the spinor-helicity basis [142].
 So the rest of the amplitude should produce something like $m^{2-2S}\langle{\bf 1}{\bf 3}\rangle^{2S-2}$.
@@
 As the term in square brackets in Eq.
-(18) is roughly the QED Compton amplitude for matter of spin one, this should be proportional to
+(20) is roughly the QED Compton amplitude for matter of spin one, in the spinor-helicity basis this should be proportional to
 $\frac{\langle 4|{\bf 1}|2\bigr{]}^{2}}{m^{4}t}\big{|}\langle{\bf 1}|P|{\bf 3}\bigr{]}\big{|}^{2}$
@@
 Again apart from the positive numerical term inside the brackets, this term reproduces the correct s-channel factorization for the amplitude with opposite helicities.
-Let us quote our proposal for the electromagnetic Compton Scattering amplitude with opposite helicities.
+Let us state our proposal for the electromagnetic Compton scattering amplitude.
+Consider first the amplitude with opposite helicities.
 Denoting by a bar this particular contribution we find, for a massive particle with spin $s=n>0$ integer:
@@
 where
-$\mathbb{A}_{n}({\bf 1},2^{+},{\bf 3},4^{-})$$\equiv$ ${\cal N}n^{2}(2n)^{n-1}\gamma(s\alpha^{\prime},n-1)\gamma(u\alpha^{\prime},n-1)$ (23)
-$\times$ $\sum_{k=0}^{n-1}\frac{(-1)^{n-1}}{\Gamma(k+1)^{2}\Gamma(-k+n)^{2}}\left(\frac{\bigl{(}q_{2}\cdot\epsilon({\bf 3})\,q_{2}\cdot\epsilon({\bf 1})\bigr{)}}{m^{2}}\right)^{n-1-k}\left(\frac{\bigl{(}q_{4}\cdot\epsilon({\bf 1})\,q_{4}\cdot\epsilon({\bf 3})\bigr{)}}{m^{2}}\right)^{k}$
+$\mathbb{A}_{n}({\bf 1},2^{+},{\bf 3},4^{-})$$\equiv$ ${\cal N}\frac{n^{2}(2n)^{n-1}(-1)^{n-1}}{\Gamma(n)^{2}}\gamma(s\alpha^{\prime},n-1)\gamma(u\alpha^{\prime},n-1)$ (25)
+$\times$ $\left(\frac{\bigl{(}q_{2}\cdot\epsilon({\bf 3})\,q_{2}\cdot\epsilon({\bf 1})\bigr{)}}{m^{2}}\right)^{n-1}\,_{2}F_{1}\left(1-n,1-n;1;\frac{\bigl{(}q_{4}\cdot\epsilon({\bf 1})\,q_{4}\cdot\epsilon({\bf 3})\bigr{)}}{\bigl{(}q_{2}\cdot\epsilon({\bf 3})\,q_{2}\cdot\epsilon({\bf 1})\bigr{)}}\right)$
 $\times$ $\epsilon^{+}_{\alpha}(2)\epsilon^{-}_{\beta}(4)\Biggl{\{}-(u-m^{2})\left[\Bigl{(}-\eta^{a_{s}c_{s}}k_{3}^{\beta}+\eta^{a_{s}\beta}q_{4}^{c_{s}}-\eta^{c_{s}\beta}q_{4}^{a_{s}}\Bigr{)}\Bigl{(}k_{3}^{\alpha}+q_{4}^{\alpha}\Bigr{)}\right.$
@@
 $+$ $\left.\eta^{a_{s}\beta}\Bigl{(}(q_{2}\cdot q_{4})\eta^{\alpha c_{s}}-q_{2}^{c_{s}}q_{4}^{\alpha}\Bigr{)}+k_{3}^{\alpha}\left(q_{4}^{c_{s}}\eta^{a_{s}\beta}-q_{4}^{a_{s}}\eta^{c_{s}\beta}\right)\right]$
-$+$ $\frac{1}{2}\eta^{\alpha\beta}\eta^{a_{s}c_{s}}(s-m^{2})(u-m^{2})\Biggr{\}}\epsilon_{a_{s}}({\bf 1})\epsilon_{c_{s}}({\bf 3}).$
+$+$ $\frac{1}{2}\eta^{\alpha\beta}\eta^{a_{s}c_{s}}(s-m^{2})(u-m^{2})\Biggr{\}}\epsilon_{a_{s}}({\bf 1})\epsilon_{c_{s}}({\bf 3})$
+where we used that
+$\sum_{k=0}^{n-1}\frac{B^{k}A^{-k+n-1}}{\Gamma(k+1)^{2}\Gamma(n-k)^{2}}=\frac{A^{n-1}\,_{2}F_{1}\left(1-n,1-n;1;\frac{B}{A}\right)}{\Gamma(n)^{2}}$
+and $\,{}_{2}F_{1}\left(a,b;c;x\right)$ is the hypergeometric function.
 Notice that we have added an arbitrary normalization factor ${\cal N}(n)$ (with ${\cal N}(1)=1$) in order to take care of the multiplicative numerical factor that appears in the factorization of the physical channels.
-Let us summarize the three constraints we have imposed:
+That is,
+${\cal N}(n)=\left(\frac{4^{n-1}n^{n+1}\Gamma\left(n-\frac{1}{2}\right)[\gamma(n,n-1)]^{2}}{\sqrt{\pi}\Gamma(n)^{3}}\right)^{-1}.$
+As for the all plus amplitude, as discussed we obtain the usual result (up to an irrelevant positive constant) which we quote here for completeness:
+$\bar{{\cal A}}_{n}({\bf 1},2^{+},{\bf 3},4^{+})=2ie^{2}m^{2-2n}\frac{\bigl{[}24\bigr{]}^{2}\langle{\bf 1}{\bf 3}\rangle^{2n}}{(s-m^{2})(u-m^{2})}\left(\frac{n^{2}[\gamma(n,n-1)]^{2}}{\Gamma(n+1)}\right).$ (26)
+[…]
+In summary, our claim is: Open string-theory amplitudes can potentially contain amplitudes that describe compact astrophysical bodies as specific contributions (among many others) in disguise inside the full amplitude  [^7] Indeed, as argued in Ref.
+[105], 3-point string amplitudes produce couplings which are “maximally complex” in that it represents the most general form of the three-point amplitude for one massless and two equal mass legs with spin $S$ but with the corresponding couplings assuming definite values.
+This amplitude includes, of course, the minimal coupling amplitude which describes root-Kerr in the classical limit..
+To extract this hidden contribution to the amplitude, all other contributions must be projected out, and we propose that this can be achieved by imposing the following three constraints:
   • Soft momentum transfer, $q^{2}\lll 1/\alpha^{\prime}$;
@@
   • Consistent factorization in the physical channels for any helicity choice.
-[…]
-$\epsilon^{\mu}({\bf 3})\epsilon^{\nu}({\bf 1})\to\bar{\epsilon}^{\mu}({\bf 1})\epsilon^{\nu}({\bf 1})=-\langle S^{(\mu}S^{\nu)}\rangle+\frac{i}{2m}\varepsilon^{\mu\nu\rho\sigma}k_{1\rho}\langle S_{\sigma}\rangle-\Theta^{\mu\nu}\equiv\Pi^{\mu\nu}$ (24)
-up to subleading terms in the classical limit, where $\Theta^{\mu\nu}$ is the spin-1 projector. For the spin-S case we take appropriate tensor products:
-$\bar{\epsilon}^{\mu_{1}}({\bf 1})\cdots\bar{\epsilon}^{\mu_{S}}({\bf 1})\epsilon^{\nu_{1}}({\bf 1})\cdots\epsilon^{\nu_{S}}({\bf 1})=\Pi^{\mu_{1}\nu_{1}}\otimes\Pi^{\mu_{2}\nu_{2}}\otimes\cdots\otimes\Pi^{\mu_{S}\nu_{S}}$ (25)
-where a symmetrised expectation value is understood over products of the spin operator.
-Now one inserts Eq.
-(25) in Eq.
-(23) and takes $n\to\infty$.
-In this limit one may identify $\langle(q\cdot S/m)^{k}\rangle$ with the classical power law $(q\cdot a)^{k}$, where $a^{\mu}$ is the classical ring-radius.
-We define the resulting expression as being the classical amplitude associated with (23).
-It is easy to see that, in the classical limit, the summation in (23) will generate powers such as $(q_{2}\cdot a)^{2l}(q_{4}\cdot a)^{2p}$, $l,p$ being integers smaller then $n$, times a piece that is proportional to the spin-1 case.
-This multiplicative piece satisfies the so-called spin-shift symmetry [22, 25, 43], $a_{i}^{\mu}\to a_{i}^{\mu}+\iota_{i}q^{\mu}/q^{2}$, where $q$ is the momentum transfer and $\iota_{i}$ is an arbitrary parameter.
+Let us see how such a reasoning can also be used to extract the correct 3-point root-Kerr amplitude from the corresponding superstring amplitude.
+The relevant three-point amplitude is given by [145, 146]
+$A_{3}({\bf 1},{\bf 2},q)$ $=$ $e(2\alpha^{\prime})^{S}(S-1)!\sum_{j=0}^{S}\frac{(-\epsilon({\bf 1})\cdot\epsilon({\bf 2}))^{j}}{(2\alpha^{\prime})^{j}j![(S-j)!]^{2}}$
+$\times$ $\left(-j(\epsilon(q)\cdot k_{1})(-\epsilon({\bf 1})\cdot q\,\epsilon({\bf 2})\cdot q)^{S-j}+\frac{S(S-j)}{2\alpha^{\prime}}\epsilon_{\mu}({\bf 2})f^{\mu\nu}\epsilon_{\nu}({\bf 1})(-\epsilon({\bf 1})\cdot q\,\epsilon({\bf 2})\cdot q)^{S-j-1}\right)$
+where, as above, $f^{\mu\nu}=2q^{[\mu}\epsilon^{\nu]}(q)$.
+The field-theory limit proposed here requires a soft expansion and, as argued before, this is parametrized by $\alpha^{\prime}$.
+Moreover, we also consider the compactification mentioned above for the massive polarizations.
+This effectively implies only the term $j=S$ in the above sum is kept, which leads us to
+$A_{3}({\bf 1},{\bf 2},q)=\frac{e}{\sqrt{2}}(-1)^{S}mx_{1}(\epsilon({\bf 1})\cdot\epsilon({\bf 2}))^{S}$ (28)
+where the $x_{1}$ factor is defined as usual by $x_{1}=\sqrt{2}(k_{1}\cdot\epsilon(q))/m$. In the classical limit, $x_{1}$ is $m$-independent [113]. On the other hand, we can use that
+$(\epsilon({\bf 1})\cdot\epsilon({\bf 2}))^{S}=\frac{1}{m^{2S}}\langle{\bf 2}{\bf 1}\rangle^{S}\bigl{[}{\bf 1}{\bf 2}\bigr{]}^{S}$ (29)
+and, from three-point special kinematics:
+$\bigl{[}{\bf 1}{\bf 2}\bigr{]}=\langle{\bf 1}{\bf 2}\rangle-\frac{1}{mx_{1}}\langle{\bf 1}q\rangle\langle q{\bf 2}\rangle$ (30)
+so we can argue that
+$A_{3}({\bf 1},{\bf 2},q^{+})$ $=$ $\frac{e}{\sqrt{2}}mx_{1}^{+}\left(\frac{\langle{\bf 1}{\bf 2}\rangle}{m}\right)^{2S}\Bigl{(}1+{\cal O}(1/m)\Bigr{)}$
+$A_{3}({\bf 1},{\bf 2},q^{-})$ $=$ $\frac{e}{\sqrt{2}}mx_{1}^{-}\left(\frac{\bigl{[}{\bf 1}{\bf 2}\bigr{]}}{m}\right)^{2S}\Bigl{(}1+{\cal O}(1/m)\Bigr{)}.$ (31)
+In the end, we are able to extract from the full 3-point amplitude the minimal coupling quantum amplitudes which describe root-Kerr in the classical limit.
+3.4 Classical limit of the opposite-helicity amplitude
+Now let us study the classical limit of Eq.
+(24) ($\hbar\to 0$, $S\to\infty$, $\hbar s$ fixed) as this is the helicity set up that contributes to the 2PM classical amplitude.
+As mentioned above, there are many works designed to tackle the problem of providing an ansatz for a local Compton amplitude.
+For practical terms, we choose here to compare our results with the ones derived in Ref.
+[106] as their presentation is more suitable for our analysis.
+Throughout this subsection, we will employ many of their definitions and manipulations to facilitate comparison.
+Many of the details which we will not display here for the sake of brevity and clarity can be found in such a reference and we refer the interested reader to it.
+One may start by rewriting Eq. (25) as
+$\mathbb{A}_{n}({\bf 1},2^{+},{\bf 3},4^{-})$ $=$ $-{\cal N}\frac{n^{2}(2n)^{n-1}(-1)^{n-1}}{\Gamma(n)^{2}}\gamma(s\alpha^{\prime},n-1)\gamma(u\alpha^{\prime},n-1)\Bigl{(}\langle{\bf 3}4\rangle\bigl{[}{\bf 1}2\bigr{]}+\langle{\bf 1}4\rangle\bigl{[}{\bf 3}2\bigr{]}\Bigr{)}^{2}$ (32)
+$\times$ $\left(\frac{\bigl{(}(q_{2}\cdot\rho)^{2}-(q_{2}\cdot\bar{\rho})^{2}\bigr{)}}{2m^{4}}\right)^{n-1}\,_{2}F_{1}\left(1-n,1-n;1;\frac{\bigl{(}(q_{4}\cdot\rho)^{2}-(q_{4}\cdot\bar{\rho})^{2}\bigr{)}}{\bigl{(}(q_{2}\cdot\rho)^{2}-(q_{2}\cdot\bar{\rho})^{2}\bigr{)}}\right)$
+where we wrote the product of massive polarizations in terms of the following two complex vectors [106]
+$\rho^{\mu}$$\equiv$ $\frac{1}{2}\Big{(}\langle\bf{3}|\sigma^{\mu}|\bf{1}]+\langle\bf{1}|\sigma^{\mu}|\bf{3}]\Big{)}\,,$
+$\bar{\rho}^{\mu}$$\equiv$ $\frac{1}{2}\Big{(}\langle\bf{3}|\sigma^{\mu}|\bf{1}]-\langle\bf{1}|\sigma^{\mu}|\bf{3}]\Big{)}\,.$ (33)
+The amplitude should display exchange symmetry between massless particles, which is not manifest here, but it is definitely inherited from the full string amplitude, as one can easily see from the above results and the appendix – this follows from the usual symmetry contained in the binomial theorem,
+$(x+y)^{n}=\sum_{k=0}^{n}{n\choose k}x^{n-k}y^{k}=\sum_{k=0}^{n}{n\choose k}x^{k}y^{n-k}.$
+We should combine this exchange symmetry with parity. Furthermore, by introducing the quantities [106]
+$W_{\pm}$$\equiv$ $\frac{m}{2}\Bigl{(}\langle{\bf 1}{\bf 3}\rangle\pm[{\bf 1}{\bf 3}]\Bigr{)}$
+$U$$\equiv$ $\frac{1}{2}\Bigl{(}\langle{\bf 1}|4|{\bf 3}]-\langle{\bf 3}|4|{\bf 1}]\Bigr{)}-m[{\bf 1}{\bf 3}]$
+$V$$\equiv$ $\frac{1}{2}\Bigl{(}\langle{\bf 1}|4|{\bf 3}]+\langle{\bf 3}|4|{\bf 1}]\Bigr{)}$
+$\chi^{\mu}$$\equiv$ $\langle 4|\sigma^{\mu}|2]$ (34)
+we can rewrite the amplitude as
+$\mathbb{A}_{n}({\bf 1},2^{+},{\bf 3},4^{-})$ $=$ $-{\cal N}\frac{n^{2}(2n)^{n-1}(-1)^{n-1}}{2\,\Gamma(n)^{2}}\gamma(s\alpha^{\prime},n-1)\gamma(u\alpha^{\prime},n-1)(\rho\cdot\chi)^{2}$ (35)
+$\times$ $\left[\left(\frac{V^{2}-\Bigl{(}4W_{-}-(U+m[{\bf 1}{\bf 3}])\Bigr{)}^{2}}{2m^{4}}\right)^{n-1}\right.$
+$\times$ $\left.\,{}_{2}F_{1}\left(1-n,1-n;1;\frac{\Bigl{(}V^{2}-(U+m[{\bf 1}{\bf 3}])^{2}\Bigr{)}}{V^{2}-\Bigl{(}4W_{-}-(U+m[{\bf 1}{\bf 3}])\Bigr{)}^{2}}\right)\right.$
+$+$ $\left.\left(\frac{V^{2}-\Bigl{(}4W_{-}+U+m\langle{\bf 1}{\bf 3}\rangle\Bigr{)}^{2}}{2m^{4}}\right)^{n-1}\right.$
+$\times$ $\left.\,{}_{2}F_{1}\left(1-n,1-n;1;\frac{\Bigl{(}V^{2}-(U+m\langle{\bf 1}{\bf 3}\rangle)^{2}\Bigr{)}}{V^{2}-\Bigl{(}4W_{-}+U+m\langle{\bf 1}{\bf 3}\rangle\Bigr{)}^{2}}\right)\right]$
+To obtain the classical amplitude, we adopt the approach based on coherent spin states, see discussion in Refs.
+[69, 106].
+The idea here is to consider the scattering of coherent spin states in place of the massive spin particles.
+The classical amplitude is the $\hbar\to 0$ limit of an infinite sum over diagonal finite-spin amplitudes
+$A_{\textrm{cl}}({\bf 1},2^{+},{\bf 3},4^{-})=\lim_{\hbar\to 0}e^{-|{\cal Z}|^{2}}\sum_{S=0}^{\infty}\frac{1}{(2S)!}\bar{{\cal A}}_{S}({\bf 1},2^{+},{\bf 3},4^{-})$ (36)
+where non-diagonal contributions are neglected and the $\hbar\to 0$ limit is taken after resumming the coherent states.
+A thorough discussion of the various $\hbar$ scalings of spin variables in the coherent-state approach can be found in Ref.
+[106].
+In turn, $|{\cal Z}|^{2}$ is the squared norm of the physical SU(2) wavefunction for a massive particle, described by SU(2) little-group spinors ${\cal Z}_{a}$ and $\bar{{\cal Z}}_{a}$ ($i=1,3$, $a=1,2$)
+$|{\bf i}\rangle$ $=$ $|i^{a}\rangle{\cal Z}_{ia}\,,\qquad\quad|\bar{\bf i}\rangle=|i^{a}\rangle\bar{{\cal Z}}_{ia}\,,$
+$|{\bf i}]$ $=$ $|i^{a}]{\cal Z}_{ia}\,,\qquad\,\quad|\bar{\bf i}]=|i^{a}]\bar{{\cal Z}}_{ia}\,,$ (37)
+where the unbarred spinors are incoming, and the barred spinors are outgoing.
+In this case $|{\cal Z}|^{2}={\cal Z}_{a}\bar{{\cal Z}}^{a}$ for each spinning particle, where ${\cal Z}^{a}=\epsilon^{ab}{\cal Z}_{b}$, $\epsilon^{ab}$ being the SU(2) Levi-Civita symbol.
+When using quantum amplitudes to calculate the classical potential, we know that classical contributions arise only from the region where internal matter lines are close to on shell [152].
+Using this reasoning, one can argue that, in the limit $\hbar\to 0$, we should have that $\gamma(s\alpha^{\prime},n-1)\gamma(u\alpha^{\prime},n-1)\approx[\gamma(n,n-1)]^{2}$.
+This result can also be obtained with the help of a soft-expansion argument.
+Moreover, summing over all possible spin states in our case means that the sum should start at $n=1$.
+Hence we claim that the classical amplitude is obtained with
+$A_{\textrm{cl}}({\bf 1},2^{+},{\bf 3},4^{-})$ $=$ $-\lim_{\hbar\to 0}\frac{ie^{2}}{\left(s-m^{2}\right)\left(u-m^{2}\right)}e^{-|{\cal Z}|^{2}}(\rho\cdot\chi)^{2}$ (38)
+$\times$ $\sum_{n=1}^{\infty}\frac{1}{(2(n-1))!}\left(\frac{\sqrt{\pi}\Gamma(n)}{4^{n-1}\Gamma\left(n-\frac{1}{2}\right)}\right)(-1)^{n-1}\left[\left(\frac{V^{2}-\Bigl{(}4W_{-}-(U+m[{\bf 1}{\bf 3}])\Bigr{)}^{2}}{2m^{4}}\right)^{n-1}\right.$
+$\times$ $\left.\,{}_{2}F_{1}\left(1-n,1-n;1;\frac{\Bigl{(}V^{2}-(U+m[{\bf 1}{\bf 3}])^{2}\Bigr{)}}{V^{2}-\Bigl{(}4W_{-}-(U+m[{\bf 1}{\bf 3}])\Bigr{)}^{2}}\right)\right.$
+$+$ $\left.\left(\frac{V^{2}-\Bigl{(}4W_{-}+U+m\langle{\bf 1}{\bf 3}\rangle\Bigr{)}^{2}}{2m^{4}}\right)^{n-1}\right.$
+$\times$ $\left.\,{}_{2}F_{1}\left(1-n,1-n;1;\frac{\Bigl{(}V^{2}-(U+m\langle{\bf 1}{\bf 3}\rangle)^{2}\Bigr{)}}{V^{2}-\Bigl{(}4W_{-}+U+m\langle{\bf 1}{\bf 3}\rangle\Bigr{)}^{2}}\right)\right].$
+To deal with the term $e^{-|{\cal Z}|^{2}}(\rho\cdot\chi)^{2}$, let us come back to the well established expression (12), which is valid for $S\leq 1$; its classical limit is given by
+$A_{\textrm{cl}}({\bf 1},2^{+},{\bf 3},4^{-})$ $=$ $-\lim_{\hbar\to 0}\frac{ie^{2}\,(k_{1}\cdot\chi)^{2}}{(s-m^{2})(u-m^{2})}e^{-|{\cal Z}|^{2}}\sum_{S=0}^{\infty}\frac{1}{(2S)!}\left(\frac{\rho\cdot\chi}{k_{1}\cdot\chi}\right)^{2S}+{\cal O}(a^{3})$ (39)
+$=$ $-\lim_{\hbar\to 0}\frac{ie^{2}\,(k_{1}\cdot\chi)^{2}}{(s-m^{2})(u-m^{2})}e^{-|{\cal Z}|^{2}+\frac{\rho\cdot\chi}{k_{1}\cdot\chi}}+{\cal O}(a^{3})$
+where the transverse spacelike vector $a^{\mu}$, with $k_{1}\cdot a=0$, is called the ring radius whose magnitude $|a|=\sqrt{a_{\mu}a^{\mu}}$ in the case of a Kerr black hole would correspond to the size of the ring singularity.
+Next, let us introduce four spin-dependent, helicity-independent and dimensionless variables [106]:
+$x$ $=$ $a\cdot(q_{2}-q_{4})\,,\qquad\;\,\qquad y=a\cdot(q_{2}+q_{4})\,,$
+$z$ $=$ $|a|\frac{k_{1}\cdot(q_{2}-q_{4})}{m}\,,\qquad\quad w=\frac{(a\cdot\chi)\;[k_{1}\cdot(q_{2}-q_{4})]}{k_{1}\cdot\chi}\,.$ (40)
+Using that $|{\cal Z}|^{2}=2m|a|$ and the classical limit of $\rho\cdot\chi$ discussed in Ref. [106], we find that
+$\lim_{\hbar\to 0}e^{-|{\cal Z}|^{2}+\frac{\rho\cdot\chi}{k_{1}\cdot\chi}}=e^{x-w}.$ (41)
+Therefore, using the fact that $k_{1}\cdot(q_{2}+q_{4})\sim\hbar^{2}$ and hence we can take $[k_{1}\cdot(q_{2}+q_{4})]^{2}\Big{|}_{\hbar\to 0}\approx 0$, we finally obtain the classical amplitude valid up to $a^{3}$:
+$A_{\textrm{cl}}({\bf 1},2^{+},{\bf 3},4^{-})=ie^{2}\frac{(k_{1}\cdot\chi)^{2}}{[k_{1}\cdot(q_{2}-q_{4})]^{2}}e^{x}\left(1-w+\frac{w^{2}}{2}\right)+{\cal O}(a^{3}).$ (42)
+We expand this way since $w$ can only emerge at most quadratically so that it will not generate the spurious pole discussed in the literature.
+Therefore, comparing with our amplitude for generic integer $S$, we learn that
+$\lim_{\hbar\to 0}e^{-|{\cal Z}|^{2}}(\rho\cdot\chi)^{2}=e^{x}\left(1-w+\frac{w^{2}}{2}\right).$ (43)
+We have to be careful when inserting this into our amplitude in order to avoid double counting the lowest order terms in the expansion in powers of $a$.
+In other words, the contribution for $n=1$ should be part of the low-orders expansion of the exponential $e^{x-w}$.
+This can be easily achieved by isolating the $n=1$ term from the rest and taking due care of the expansion of $e^{x-w}$ so that we have at most quadratic orders in $w$ and absence of the aforementioned double counting.
+In this way, we obtain
+$A_{\textrm{cl}}({\bf 1},2^{+},{\bf 3},4^{-})$ $=$ $ie^{2}\frac{(k_{1}\cdot\chi)^{2}}{[k_{1}\cdot(q_{2}-q_{4})]^{2}}\Bigl{[}1+(e^{x}-1)\Bigr{]}\left[1+\left(-w+\frac{w^{2}}{2}\right)\right]$ (44)
+$+$ $ie^{2}\frac{(k_{1}\cdot\chi)^{2}}{[k_{1}\cdot(q_{2}-q_{4})]^{2}}\lim_{\hbar\to 0}\sum_{n=2}^{\infty}\frac{1}{(2(n-1))!}\left(\frac{\sqrt{\pi}\Gamma(n)}{4^{n-1}\Gamma\left(n-\frac{1}{2}\right)}\right)(-1)^{n-1}$
+$\times$ $\left\{\left(\frac{w^{2}}{2}-w\right)(e^{x}-1)\left[\left(\frac{V^{2}-\Bigl{(}4W_{-}-(U+m[{\bf 1}{\bf 3}])\Bigr{)}^{2}}{2m^{4}}\right)^{n-1}\right.\right.$
+$\times$ $\left.\left.\,{}_{2}F_{1}\left(1-n,1-n;1;\frac{\Bigl{(}V^{2}-(U+m[{\bf 1}{\bf 3}])^{2}\Bigr{)}}{V^{2}-\Bigl{(}4W_{-}-(U+m[{\bf 1}{\bf 3}])\Bigr{)}^{2}}\right)\right.\right.$
+$+$ $\left.\left.\left(\frac{V^{2}-\Bigl{(}4W_{-}+U+m\langle{\bf 1}{\bf 3}\rangle\Bigr{)}^{2}}{2m^{4}}\right)^{n-1}\right.\right.$
+$\times$ $\left.\left.\,{}_{2}F_{1}\left(1-n,1-n;1;\frac{\Bigl{(}V^{2}-(U+m\langle{\bf 1}{\bf 3}\rangle)^{2}\Bigr{)}}{V^{2}-\Bigl{(}4W_{-}+U+m\langle{\bf 1}{\bf 3}\rangle\Bigr{)}^{2}}\right)\right]\right.$
+$+$ $\left.\left(\frac{w^{2}}{2}-w\right)\left[\left(\frac{V^{2}-\Bigl{(}4W_{-}-(U+m[{\bf 1}{\bf 3}])\Bigr{)}^{2}}{2m^{4}}\right)^{n-1}\right.\right.$
+$\times$ $\left.\left.\,{}_{2}F_{1}\left(1-n,1-n;1;\frac{\Bigl{(}V^{2}-(U+m[{\bf 1}{\bf 3}])^{2}\Bigr{)}}{V^{2}-\Bigl{(}4W_{-}-(U+m[{\bf 1}{\bf 3}])\Bigr{)}^{2}}\right)\right.\right.$
+$+$ $\left.\left.\left(\frac{V^{2}-\Bigl{(}4W_{-}+U+m\langle{\bf 1}{\bf 3}\rangle\Bigr{)}^{2}}{2m^{4}}\right)^{n-1}\right.\right.$
+$\times$ $\left.\left.\,{}_{2}F_{1}\left(1-n,1-n;1;\frac{\Bigl{(}V^{2}-(U+m\langle{\bf 1}{\bf 3}\rangle)^{2}\Bigr{)}}{V^{2}-\Bigl{(}4W_{-}+U+m\langle{\bf 1}{\bf 3}\rangle\Bigr{)}^{2}}\right)\right]\right.$
+$+$ $\left.(e^{x}-1)\left[\left(\frac{V^{2}-\Bigl{(}4W_{-}-(U+m[{\bf 1}{\bf 3}])\Bigr{)}^{2}}{2m^{4}}\right)^{n-1}\right.\right.$
+$\times$ $\left.\left.\,{}_{2}F_{1}\left(1-n,1-n;1;\frac{\Bigl{(}V^{2}-(U+m[{\bf 1}{\bf 3}])^{2}\Bigr{)}}{V^{2}-\Bigl{(}4W_{-}-(U+m[{\bf 1}{\bf 3}])\Bigr{)}^{2}}\right)\right.\right.$
+$+$ $\left.\left.\left(\frac{V^{2}-\Bigl{(}4W_{-}+U+m\langle{\bf 1}{\bf 3}\rangle\Bigr{)}^{2}}{2m^{4}}\right)^{n-1}\right.\right.$
+$\times$ $\left.\left.\,{}_{2}F_{1}\left(1-n,1-n;1;\frac{\Bigl{(}V^{2}-(U+m\langle{\bf 1}{\bf 3}\rangle)^{2}\Bigr{)}}{V^{2}-\Bigl{(}4W_{-}+U+m\langle{\bf 1}{\bf 3}\rangle\Bigr{)}^{2}}\right)\right]\right\}.$
+The sum over the coherent states will produce an entire function which can be expanded according to the discussion in Ref.
+[106].
+Hence the final result for the classical opposite-helicity amplitude reads
+$A_{\textrm{cl}}({\bf 1},2^{+},{\bf 3},4^{-})$ $=$ $ie^{2}\frac{(k_{1}\cdot\chi)^{2}}{[k_{1}\cdot(q_{2}-q_{4})]^{2}}\Bigl{[}1+(e^{x}-1)\Bigr{]}\left[1+\left(-w+\frac{w^{2}}{2}\right)\right]$ (45)
+$+$ $ie^{2}\frac{(k_{1}\cdot\chi)^{2}}{[k_{1}\cdot(q_{2}-q_{4})]^{2}}\sum_{n=2}^{\infty}\frac{1}{(2(n-1))!}\left(\frac{\sqrt{\pi}\,\Gamma(n)}{4^{n-1}\Gamma\left(n-\frac{1}{2}\right)}\right)$
+$\times$ $\left\{(e^{x}-1)\left(\frac{w^{2}}{2}-w\right)\left[\Bigl{(}({x}+5{y})^{2}-{z}^{2}\Bigr{)}^{n-1}\,_{2}F_{1}\left(1-n,1-n;1;\frac{({x}+{y})^{2}-{z}^{2}}{({x}+5{y})^{2}-{z}^{2}}\right)\right.\right.$
+$+$ $\left.\left.\Bigl{(}({x}-5{y})^{2}-{z}^{2}\Bigr{)}^{n-1}\,_{2}F_{1}\left(1-n,1-n;1;\frac{({x}-{y})^{2}-{z}^{2}}{({x}-5{y})^{2}-{z}^{2}}\right)\right]\right.$
+$+$ $\left.\left(\frac{w^{2}}{2}-w\right)\left[\Bigl{(}({x}+5{y})^{2}-{z}^{2}\Bigr{)}^{n-1}\,_{2}F_{1}\left(1-n,1-n;1;\frac{({x}+{y})^{2}-{z}^{2}}{({x}+5{y})^{2}-{z}^{2}}\right)\right.\right.$
+$+$ $\left.\left.\Bigl{(}({x}-5{y})^{2}-{z}^{2}\Bigr{)}^{n-1}\,_{2}F_{1}\left(1-n,1-n;1;\frac{({x}-{y})^{2}-{z}^{2}}{({x}-5{y})^{2}-{z}^{2}}\right)\right]\right.$
+$+$ $\left.(e^{x}-1)\left[\Bigl{(}({x}+5{y})^{2}-{z}^{2}\Bigr{)}^{n-1}\,_{2}F_{1}\left(1-n,1-n;1;\frac{({x}+{y})^{2}-{z}^{2}}{({x}+5{y})^{2}-{z}^{2}}\right)\right.\right.$
+$+$ $\left.\left.\Bigl{(}({x}-5{y})^{2}-{z}^{2}\Bigr{)}^{n-1}\,_{2}F_{1}\left(1-n,1-n;1;\frac{({x}-{y})^{2}-{z}^{2}}{({x}-5{y})^{2}-{z}^{2}}\right)\right]\right\}+{\cal O}(\hbar).$
+Even though the appearance of the hypergeometric function makes the result sound a little sophisticated, it can be easily shown that the function actually reduces to a polynomial.
+Hence the final result is completely expressed in terms of entire functions.
+Expanding the result up to ${\cal O}(a^{3})$, our results agree with those in Ref.
+[106].
+At higher orders, our results seem to differ.
+Amusingly, setting $y=0$, we find the same operators in the expansion in powers of $a$ (i.e.
+powers of $x$ and $z$), albeit accompanied by different coefficients (at least up to ${\cal O}(x^{6})$ and ${\cal O}(z^{6})$).
+On the other hand, the $n=1$ case satisfies the so-called spin-shift symmetry [100, 102, 75], $a_{i}^{\mu}\to a_{i}^{\mu}+\iota_{i}q^{\mu}/q^{2}$, where $q$ is the momentum transfer and $\iota_{i}$ is an arbitrary parameter.
 As emphasized by Ref.
@@
 We conclude that the classical limit of our proposal for the opposite-helicity amplitude should fail to obey this symmetry for $n>1$.
-However, the massive-spin one ($n=1$ case) opposite-helicity Compton amplitude respects the spin-shift symmetry, as it should.
 4 Gravitational Compton amplitudes for all spins
+4.1 Double copy and first mass level
 Now let us study the gravity Compton amplitude. Using our previous prescription for deriving the associated field-theory amplitudes, we get
-$M^{\alpha\gamma,\beta\delta}_{4}({\bf 1},2,{\bf 3},4)=8i\left(\frac{(-1)^{n}}{n[\Gamma(n)]^{4}}\right)\kappa^{2}m^{2}\frac{1}{\alpha^{\prime}t(s\alpha^{\prime}-4n)(u\alpha^{\prime}-4n)}\left[{\cal A}_{0}^{\alpha\beta}{\cal A}_{0}^{\gamma\delta}\right]_{\alpha^{\prime}\to\alpha^{\prime}/4}$ (26)
-where we have expanded $\sin\left(\pi x\right)$ around $x=n$ and also restored the gravity coupling constants, $M_{N}\to\left(\kappa/2\right)^{N-2}M_{N}$ [76], where $\kappa^{2}=32\pi G$.
+$M^{\alpha\gamma,\beta\delta}_{4}({\bf 1},2,{\bf 3},4)=64i\left(\frac{(-1)^{n}}{n[\Gamma(n)]^{4}}\right)\kappa^{2}m^{2}\frac{1}{\alpha^{\prime}t(s\alpha^{\prime}-4n)(u\alpha^{\prime}-4n)}\left[{\cal A}_{0}^{\alpha\beta}{\cal A}_{0}^{\gamma\delta}\right]_{\alpha^{\prime}\to\alpha^{\prime}/4}$ (46)
+where we have expanded $\sin\left(\pi x\right)$ around $x=n$ and also restored the gravity coupling constants, $M_{N}\to\kappa^{N-2}M_{N}$ [3] ($N$ is the number of external particles), where $\kappa^{2}=32\pi G$.
 The first mass level $n=1$ corresponds in our prescription to a massive spin-2 particle since here we are compactifying two polarizations of each of the gauge amplitudes.
@@
 [151].
-In fact, up to quartic order in classical spin our results also match with the ones given in Ref.
-[22].
+In fact, up to fifth order in classical spin our result for the opposite helicity amplitude also matches with the one given in Ref.
+[100], as we will show below.
+For opposite helicities, we can write that
+$M_{S}({\bf 1},2^{++},{\bf 3},4^{--})$ $=$ $-i\frac{\kappa^{2}}{4\,m^{2S}}\frac{\langle 4|{\bf 1}|2\bigr{]}^{4}}{(s-m^{2})(u-m^{2})t}\langle{\bf 3}|^{2S}\exp\left(i\frac{q_{4}^{\mu}\bar{\epsilon}_{4}^{\nu}{\cal J}_{1\mu\nu}}{\bar{p}\cdot\bar{\epsilon}_{4}}\right)|{\bf 1}\rangle^{2S}$ (47)
+$=$ $-i\frac{\kappa^{2}}{4}\frac{1}{(s-m^{2})(u-m^{2})t}\langle 4|{\bf 1}|2\bigl{]}^{4-2S}\Bigl{(}\langle{\bf 3}4\rangle\bigl{[}{\bf 1}2\bigr{]}+\langle{\bf 1}4\rangle\bigl{[}{\bf 3}2\bigr{]}\Bigr{)}^{2S}\,\,\,\,\,(S\leq 2)$
+which is just the standard result [113, 142].
+As above, one can also perform the same analysis for the same-helicity amplitude, finding the usual value quoted in the literature (see the general expression below).
+4.2 Double copy and all mass levels
+Let us now present the gravitational Compton amplitude for even positive integer $S$.
 After imposing the three constraints listed above, the KLT relations allows us to derive our candidate opposite helicity gravity Compton amplitude which is valid for all quantum spins $S=2n>0$:
-¯$\bar{\cal M}_{S}({\bf 1},2^{++},{\bf 3},4^{--})$ $=$ $8i\left(\frac{(-1)^{n}}{n[\Gamma(n)]^{4}}\right)\kappa^{2}m^{2}\frac{1}{\alpha^{\prime}t(\alpha^{\prime}s-4n)(\alpha^{\prime}u-4n)}$ (27)
+¯$\bar{\cal M}_{S}({\bf 1},2^{++},{\bf 3},4^{--})$ $=$ $64i\left(\frac{(-1)^{n}}{n[\Gamma(n)]^{4}}\right)\kappa^{2}m^{2}\frac{1}{\alpha^{\prime}t(\alpha^{\prime}s-4n)(\alpha^{\prime}u-4n)}$ (48)
 $\times$ $\left[\alpha^{\prime 4}\mathbb{A}_{S/2}({\bf 1},2^{+},{\bf 3},4^{-})\mathbb{A}_{S/2}({\bf 1},2^{+},{\bf 3},4^{-})\right]_{\alpha^{\prime}\to\alpha^{\prime}/4}.$
-One can also similarly construct the all plus gravitational amplitude from the corresponding electromagnetic case and the result matches the one known in the literature [69], apart from a positive multiplicative numerical factor.
-Moreover, the classical limit of the opposite-helicity amplitude does not satisfy the spin-shift symmetry for quantum spins greater than 2; in any case, the massive-spin 2 opposite-helicity amplitude ($S=2$) obtained from the above double copy formula preserves this symmetry.
-We impose the constraints on the gravitational amplitude after performing the double copy procedure on the full string amplitude.
+One can also similarly construct the all plus gravitational amplitude from the corresponding electromagnetic case and the result matches the one known in the literature [142]
+¯$\bar{\cal M}_{S}({\bf 1},2^{++},{\bf 3},4^{++})=-im^{4-2S}\frac{\kappa^{2}}{4}\frac{\bigl{[}24\bigr{]}^{4}}{t}\frac{\langle{\bf 1}{\bf 3}\rangle^{2S}}{[s-m^{2}][u-m^{2}]},$ (49)
+apart from a positive multiplicative numerical factor (not shown in the expression above). This formula is valid to all $S$.
+Let us see how the 3-point Kerr amplitude can also be identified as a contribution in the full closed string amplitude, after imposing the above constraints.
+We simply repeat the calculation for the closed superstring exploiting the simple KLT relation at three points
+$M_{3}({\bf 1}_{S},{\bf 2}_{S},k)=i\kappa\left[\frac{1}{e}A_{3}({\bf 1}_{S/2},{\bf 2}_{S/2},k)\right]\left[\frac{1}{e}A_{3}({\bf 1}_{S/2},{\bf 2}_{S/2},k)\right],$ (50)
+where as above we perform the rescaling $\alpha^{\prime}\to\alpha^{\prime}/4$ which is counterbalanced by the fact that the leading Regge closed string states produce $\alpha^{\prime}=2(S-2)/m^{2}$.
+We find that
+$M_{3}({\bf 1},{\bf 2},k^{++})$ $=$ $(-1)^{S}\frac{i\kappa}{2}(mx_{1}^{+})^{2}\left(\frac{\langle{\bf 1}{\bf 2}\rangle}{m}\right)^{2S}\Bigl{(}1+{\cal O}(1/m)\Bigr{)}$
+$M_{3}({\bf 1},{\bf 2},k^{--})$ $=$ $(-1)^{S}\frac{i\kappa}{2}(mx_{1}^{-})^{2}\left(\frac{\bigl{[}{\bf 1}{\bf 2}\bigr{]}}{m}\right)^{2S}\Bigl{(}1+{\cal O}(1/m)\Bigr{)}$ (51)
+which of course is the usual result for the minimally coupled gravitational amplitude.
+Similar to the electromagnetic case, our claim here is that it is possible that closed string-theory amplitudes secretly enclose compact-body amplitudes.
+Furthermore, observe that we impose the constraints on the gravitational amplitude after performing the double copy procedure on the full string amplitude.
 This could make the double-copy construction compatible with factorization poles that match the known 3-point amplitudes.
 We believe this happens because our amplitude inherits some aspects of the full original string amplitude; it is a “stringy” quantum-field-theory amplitude.
-To confirm this it is also necessary to exploit the R sector of string theory.
+4.3 Classical limit of the opposite-helicity gravitational amplitude
+Here we will present the classical limit of Eq.
+(48).
+Again, as discussed above there are many works in the literature concerning the calculation of this amplitude, but for practical purposes we choose to use Ref.
+[107] as our main source of comparison.
+Many details that will not be exhibited here can be found in such a work.
+Again we start by rewriting the amplitude in terms of the variables $W_{\pm},U,V,\rho$ and $\chi$.
+We find that
+¯$\bar{\cal M}_{S}({\bf 1},2^{++},{\bf 3},4^{--})$ $=$ $i\left(\frac{(-1)^{n}}{[\Gamma(n)]^{4}}\right)\kappa^{2}\frac{1}{t(s-m^{2})(u-m^{2})}$ (52)
+$\times$ $\left({\cal N}\frac{n^{2}(2n)^{n-1}(-1)^{n-1}}{2\,\Gamma(n)^{2}}\gamma(s\alpha^{\prime}/4,n-1)\gamma(u\alpha^{\prime}/4,n-1)\right)^{2}(\rho\cdot\chi)^{4}$
+$\times$ $\left[\left(\frac{V^{2}-\Bigl{(}4W_{-}-(U+m[{\bf 1}{\bf 3}])\Bigr{)}^{2}}{2m^{4}}\right)^{n-1}\right.$
+$\times$ $\left.\,{}_{2}F_{1}\left(1-n,1-n;1;\frac{\Bigl{(}V^{2}-(U+m[{\bf 1}{\bf 3}])^{2}\Bigr{)}}{V^{2}-\Bigl{(}4W_{-}-(U+m[{\bf 1}{\bf 3}])\Bigr{)}^{2}}\right)\right.$
+$+$ $\left.\left(\frac{V^{2}-\Bigl{(}4W_{-}+U+m\langle{\bf 1}{\bf 3}\rangle\Bigr{)}^{2}}{2m^{4}}\right)^{n-1}\right.$
+$\times$ $\left.\,{}_{2}F_{1}\left(1-n,1-n;1;\frac{\Bigl{(}V^{2}-(U+m\langle{\bf 1}{\bf 3}\rangle)^{2}\Bigr{)}}{V^{2}-\Bigl{(}4W_{-}+U+m\langle{\bf 1}{\bf 3}\rangle\Bigr{)}^{2}}\right)\right]^{2}$
+where we recall that $S=2n$. The classical limit is obtained with
+$M_{\textrm{cl}}({\bf 1},2^{++},{\bf 3},4^{--})$ $=$ $\lim_{\hbar\to 0}\frac{i\,\kappa^{2}}{4}\frac{1}{t(s-m^{2})(u-m^{2})}e^{-|{\cal Z}|^{2}}(\rho\cdot\chi)^{4}$ (53)
+$\times$ $\sum_{n=1}^{\infty}\frac{1}{(2(n-1))!}\left(\frac{\sqrt{\pi}}{4^{n-1}\Gamma(n)\Gamma\left(n-\frac{1}{2}\right)}\right)^{2}(-1)^{3n}$
+$\times$ $\left[\left(\frac{V^{2}-\Bigl{(}4W_{-}-(U+m[{\bf 1}{\bf 3}])\Bigr{)}^{2}}{2m^{4}}\right)^{n-1}\right.$
+$\times$ $\left.\,{}_{2}F_{1}\left(1-n,1-n;1;\frac{\Bigl{(}V^{2}-(U+m[{\bf 1}{\bf 3}])^{2}\Bigr{)}}{V^{2}-\Bigl{(}4W_{-}-(U+m[{\bf 1}{\bf 3}])\Bigr{)}^{2}}\right)\right.$
+$+$ $\left.\left(\frac{V^{2}-\Bigl{(}4W_{-}+U+m\langle{\bf 1}{\bf 3}\rangle\Bigr{)}^{2}}{2m^{4}}\right)^{n-1}\right.$
+$\times$ $\left.\,{}_{2}F_{1}\left(1-n,1-n;1;\frac{\Bigl{(}V^{2}-(U+m\langle{\bf 1}{\bf 3}\rangle)^{2}\Bigr{)}}{V^{2}-\Bigl{(}4W_{-}+U+m\langle{\bf 1}{\bf 3}\rangle\Bigr{)}^{2}}\right)\right]^{2}.$
+To understand the $\hbar\to 0$ limit of the term $e^{-|{\cal Z}|^{2}}(\rho\cdot\chi)^{4}$, we again resort to the standard expression for $M_{S}({\bf 1},2^{++},{\bf 3},4^{--})$ with $S\leq 2$ given above.
+After performing similar steps as the ones carried out in the previous section, we find that
+$M_{S=2,\textrm{cl}}({\bf 1},2^{++},{\bf 3},4^{--})=-i\frac{\kappa^{2}}{4}\frac{(k_{1}\cdot\chi)^{4}}{[k_{1}\cdot(q_{2}-q_{4})]^{2}(q_{2}+q_{4})^{2}}e^{x}\left(1-w+\frac{w^{2}}{2}-\frac{w^{3}}{3!}+\frac{w^{4}}{4!}\right)+{\cal O}(a^{5}).$ (54)
+where again the expansion follows from the fact that we can have at most $w^{4}$ to prevent spurious poles from showing up.
+We can use this result in our amplitude and, as before, we prevent possible double-counting by suitably expanding $e^{x-w}$.
+In this way we obtain our final formula for the classical gravitational amplitude for opposite helicities:
+$M_{\textrm{cl}}({\bf 1},2^{++},{\bf 3},4^{--})$ $=$ $-\frac{i\,\kappa^{2}}{4}\frac{(p_{1}\cdot\chi)^{4}}{[p_{1}\cdot(q_{2}-q_{4})]^{2}(q_{2}+q_{4})^{2}}[1+(e^{x}-1)]\left[1+\left(-w+\frac{w^{2}}{2}-\frac{w^{3}}{3!}+\frac{w^{4}}{4!}\right)\right]$ (55)
+$+$ $\frac{i\,\kappa^{2}}{4}\frac{(p_{1}\cdot\chi)^{4}}{[p_{1}\cdot(q_{2}-q_{4})]^{2}(q_{2}+q_{4})^{2}}\sum_{n=2}^{\infty}\frac{1}{(2(n-1))!}\left(\frac{\sqrt{\pi}}{4^{n-1}\Gamma(n)\Gamma\left(n-\frac{1}{2}\right)}\right)^{2}(-1)^{n}$
+$\times$ $\left\{(e^{x}-1)\left(-w+\frac{w^{2}}{2}-\frac{w^{3}}{3!}+\frac{w^{4}}{4!}\right)\right.$
+$\times$ $\left.\left[\Bigl{(}({x}+5{y})^{2}-{z}^{2}\Bigr{)}^{n-1}\,_{2}F_{1}\left(1-n,1-n;1;\frac{({x}+{y})^{2}-{z}^{2}}{({x}+5{y})^{2}-{z}^{2}}\right)\right.\right.$
+$+$ $\left.\left.\Bigl{(}({x}-5{y})^{2}-{z}^{2}\Bigr{)}^{n-1}\,_{2}F_{1}\left(1-n,1-n;1;\frac{({x}-{y})^{2}-{z}^{2}}{({x}-5{y})^{2}-{z}^{2}}\right)\right]^{2}\right.$
+$+$ $\left.\left(-w+\frac{w^{2}}{2}-\frac{w^{3}}{3!}+\frac{w^{4}}{4!}\right)\left[\Bigl{(}({x}+5{y})^{2}-{z}^{2}\Bigr{)}^{n-1}\,_{2}F_{1}\left(1-n,1-n;1;\frac{({x}+{y})^{2}-{z}^{2}}{({x}+5{y})^{2}-{z}^{2}}\right)\right.\right.$
+$+$ $\left.\left.\Bigl{(}({x}-5{y})^{2}-{z}^{2}\Bigr{)}^{n-1}\,_{2}F_{1}\left(1-n,1-n;1;\frac{({x}-{y})^{2}-{z}^{2}}{({x}-5{y})^{2}-{z}^{2}}\right)\right]^{2}\right.$
+$+$ $\left.(e^{x}-1)\left[\Bigl{(}({x}+5{y})^{2}-{z}^{2}\Bigr{)}^{n-1}\,_{2}F_{1}\left(1-n,1-n;1;\frac{({x}+{y})^{2}-{z}^{2}}{({x}+5{y})^{2}-{z}^{2}}\right)\right.\right.$
+$+$ $\left.\left.\Bigl{(}({x}-5{y})^{2}-{z}^{2}\Bigr{)}^{n-1}\,_{2}F_{1}\left(1-n,1-n;1;\frac{({x}-{y})^{2}-{z}^{2}}{({x}-5{y})^{2}-{z}^{2}}\right)\right]^{2}\right\}+{\cal O}(\hbar).$
+Expanding the result up to ${\cal O}(a^{5})$, our results agree with those in Ref.
+[107].
+At higher orders, our results differ.
+Moreover, the classical limit of the opposite-helicity amplitude does not satisfy the spin-shift symmetry for quantum spins greater than 2; in any case, the massive-spin 2 opposite-helicity amplitude ($S=2$) obtained from the above double copy formula preserves this symmetry.
 5 Conclusions and perspectives
@@
 As helicity amplitudes the sum of components is generally not allowed, yet this is precisely how one can sort out the different pieces of the amplitude that unifies all the different helicity amplitudes merged into a single object.
-Remarkably we observe a similar phenomenon in the present case.
+Remarkably, we observe a similar phenomenon in the present case.
 There are other prominent features of such amplitudes which potentially justify the given interpretation. We list these here:
@@
   5. 5.
-    The gravitational amplitudes obey double copy.
-One possible conclusion is that subleading Regge trajectories seem to play an important role here.
-Indeed, as we have shown in the case of the first excited state $n=1$, the compactification down to the spin-$1$ case allowed us to recover known results in the literature and supplied the motivation to search for a valid amplitude for higher spins based on a string-theory calculation.
-Perhaps subleading Regge states could be the key to understand why this kind of compactification is needed here.
-Our results require further explorations and many of the properties listed above are calling for a better comprehension.
-For instance, our soft expansion requires some refinement, perhaps following standard procedures in the literature [78, 79, 80].
-On the other hand, we need to outline a better comparison with other works on the Compton amplitude such as Refs.
-[30, 31, 32, 33, 21, 22, 41, 58].
-For instance, one could expect that contact terms may look different as in the present case they were fixed by the string excitations.
-These comparisons are important in order to connect our results with the existing literature and also to understand better the classical limit of our candidate Compton amplitude.
-Indeed, the worldsheet interpretation for Kerr provided by Ref.
+    The gravitational quantum amplitudes obey double copy.
+As we have shown in the case of the first excited state $n=1$, the compactification down to the spin-$1$ case allowed us to recover known results in the literature and supplied the motivation to search for a valid amplitude for higher spins based on a string-theory calculation.
+We argued that this works also for the case of the 3-point amplitude.
+Nevertheless, it should be noted that this is certainly an ambiguous procedure (at best) that calls for a better comprehension.
+One possible conclusion is that subleading Regge trajectories could be the key to understanding why this kind of exotic compactification is needed here.
+While comparisons to other works could certainly be enlightening, we chose to compare our classical Compton-amplitude results to those of Refs.
+[106, 107], as we have emphasized above.
+The reason for this is that these works allow a direct parallel with our results, facilitating not only the extraction of the classical limit of the amplitudes presented here, but also making clear the physics behind the outcomes.
+Indeed, one can even speculate whether there is a UV completion within the framework of string theory for the massive higher-spin effective theory investigated in such references.
+On the other hand, our soft expansion requires some refinement, perhaps following standard procedures in the literature [153, 154, 155].
+Going further, the worldsheet interpretation for Kerr provided by Ref.
 [96] might fit well with our description.
@@
 This also motivates the study of the R sector of the superstring.
-Such studies will be reported in a separate publication elsewhere.
+We intend to explore such research directions in the future.
 Acknowledgements
@@
 with
-${\cal A}^{\alpha\beta(1)}_{4}({\bf 1},2,{\bf 3},4)$ $=$ $2i\alpha^{\prime}e^{2}(n!)^{2}\sum_{j=0}^{n}\frac{(2\alpha^{\prime})^{n-j}}{j![(n-j)!]^{2}}\eta^{a(j)c(j)}$
+${\cal A}^{\alpha\beta(1)}_{4}({\bf 1},2,{\bf 3},4)$ $=$ $2i\alpha^{\prime}e^{2}(n!)^{2}\sum_{j=0}^{n}\frac{(2\alpha^{\prime})^{n-j}}{j![(n-j)!]^{2}}\eta^{a(j)c(j)}$ (58)
 $\times$ $\sum_{k=0}^{n-j}\begin{pmatrix}n-j\\ k\end{pmatrix}(-1)^{k}q_{2}^{a(n-j-k)}q_{4}^{a(k)}\sum_{l=0}^{n-j}\begin{pmatrix}n-j\\ l\end{pmatrix}(-1)^{n-j-l}q_{2}^{c(n-j-l)}q_{4}^{c(l)}$
@@
 -- $(-s\alpha^{\prime}+k-l+n)\left(q_{4}^{a_{S}}\left(q_{2}^{\beta}\eta^{\alpha c_{S}}-q_{2}^{c_{S}}\eta^{\alpha\beta}\right)-\eta^{a_{S}\beta}\left((q_{2}\cdot q_{4})\eta^{\alpha c_{S}}-q_{2}^{c_{S}}q_{4}^{\alpha}\right)\right)\Bigr{]}$
-$\times$ $\gamma(s\alpha^{\prime},k-l+n-1)\gamma(u\alpha^{\prime},-k+l+n-1)\biggr{\}}$ (35)
-${\cal A}^{\alpha\beta(2)}_{4}({\bf 1},2,{\bf 3},4)$ $=$ $2i\alpha^{\prime}e^{2}(n!)^{2}\sum_{j=0}^{n}\frac{(2\alpha^{\prime})^{n-j}}{j![(n-j)!]^{2}}\eta^{a(j)c(j)}\left(q_{2}^{a_{S}}\eta^{\alpha c_{S}}-q_{2}^{c_{S}}\eta^{a_{S}\alpha}\right)$
+$\times$ $\gamma(s\alpha^{\prime},k-l+n-1)\gamma(u\alpha^{\prime},-k+l+n-1)\biggr{\}}$
+${\cal A}^{\alpha\beta(2)}_{4}({\bf 1},2,{\bf 3},4)$ $=$ $2i\alpha^{\prime}e^{2}(n!)^{2}\sum_{j=0}^{n}\frac{(2\alpha^{\prime})^{n-j}}{j![(n-j)!]^{2}}\eta^{a(j)c(j)}\left(q_{2}^{a_{S}}\eta^{\alpha c_{S}}-q_{2}^{c_{S}}\eta^{a_{S}\alpha}\right)$ (59)
 $\times$ $\biggl{\{}-\sum_{k=0}^{n-j}\begin{pmatrix}n-j\\ k\end{pmatrix}(-1)^{k}q_{2}^{a(n-j-k)}q_{4}^{a(k)}\sum_{l=0}^{n-j}\begin{pmatrix}n-j\\ l\end{pmatrix}(-1)^{n-j-l}q_{2}^{c(n-j-l)}q_{4}^{c(l)}$
 $\times$ $\left((-s\alpha^{\prime}+k-l+n)q_{2}^{\beta}+\alpha^{\prime}t\,k_{3}^{\beta}\right)\gamma(s\alpha^{\prime},k-l+n-1)\gamma(u\alpha^{\prime},-k+l+n-1)$
-$+$ $\frac{(n-j)\alpha^{\prime}t}{(2\alpha^{\prime})}\sum_{k=0}^{n-j-1}\begin{pmatrix}n-j-1\\ k\end{pmatrix}(-1)^{k}q_{2}^{a(n-j-1-k)}q_{4}^{a(k)}$ (43)
+$+$ $\frac{(n-j)\alpha^{\prime}t}{(2\alpha^{\prime})}\sum_{k=0}^{n-j-1}\begin{pmatrix}n-j-1\\ k\end{pmatrix}(-1)^{k}q_{2}^{a(n-j-1-k)}q_{4}^{a(k)}$
 $\times$ $\sum_{l=0}^{n-j-1}\begin{pmatrix}n-j-1\\ l\end{pmatrix}(-1)^{n-j-1-l}q_{2}^{c(n-j-1-l)}q_{4}^{c(l)}$
@@
 $+$ $\left.\eta^{\beta c_{n}}\Bigl{(}\gamma(s\alpha^{\prime},k-l+n-2)\gamma(u\alpha^{\prime},-k+l+n)q_{2}^{a_{n}}\right.$
--- $\left.\gamma(s\alpha^{\prime},k-l+n-1)\gamma(u\alpha^{\prime},-k+l+n-1)q_{4}^{a_{n}}\Bigr{)}\right]\biggr{\}}$ (47)
-${\cal A}^{\alpha\beta(3)}_{4}({\bf 1},2,{\bf 3},4)$ $=$ $2i\alpha^{\prime}e^{2}(n!)^{2}\sum_{j=0}^{n}\frac{(2\alpha^{\prime})^{n-j}}{j![(n-j)!]^{2}}\eta^{a(j)c(j)}\left(q_{4}^{c_{S}}\eta^{a_{S}\beta}-q_{4}^{a_{S}}\eta^{c_{S}\beta}\right)$
+-- $\left.\gamma(s\alpha^{\prime},k-l+n-1)\gamma(u\alpha^{\prime},-k+l+n-1)q_{4}^{a_{n}}\Bigr{)}\right]\biggr{\}}$
+${\cal A}^{\alpha\beta(3)}_{4}({\bf 1},2,{\bf 3},4)$ $=$ $2i\alpha^{\prime}e^{2}(n!)^{2}\sum_{j=0}^{n}\frac{(2\alpha^{\prime})^{n-j}}{j![(n-j)!]^{2}}\eta^{a(j)c(j)}\left(q_{4}^{c_{S}}\eta^{a_{S}\beta}-q_{4}^{a_{S}}\eta^{c_{S}\beta}\right)$ (60)
 $\times$ $\left\{\sum_{k=0}^{n-j}\begin{pmatrix}n-j\\ k\end{pmatrix}(-1)^{k}q_{2}^{a(n-j-k)}q_{4}^{a(k)}\sum_{l=0}^{n-j}\begin{pmatrix}n-j\\ l\end{pmatrix}(-1)^{n-j-l}q_{2}^{c(n-j-l)}q_{4}^{c(l)}\right.$
 $\times$ $\left.\Bigl{(}(-u\alpha^{\prime}-k+l+n)q_{4}^{\alpha}+\alpha^{\prime}t\,k_{3}^{\alpha}\Bigr{)}\gamma(s\alpha^{\prime},k-l+n-1)\gamma(u\alpha^{\prime},-k+l+n-1)\right.$
-$+$ $\left.\frac{(n-j)\alpha^{\prime}t}{(2\alpha^{\prime})}\sum_{k=0}^{n-j-1}\begin{pmatrix}n-j-1\\ k\end{pmatrix}(-1)^{k}q_{2}^{a(n-j-1-k)}q_{4}^{a(k)}\right.$ (55)
+$+$ $\left.\frac{(n-j)\alpha^{\prime}t}{(2\alpha^{\prime})}\sum_{k=0}^{n-j-1}\begin{pmatrix}n-j-1\\ k\end{pmatrix}(-1)^{k}q_{2}^{a(n-j-1-k)}q_{4}^{a(k)}\right.$
 $\times$ $\left.\sum_{l=0}^{n-j-1}\begin{pmatrix}n-j-1\\ l\end{pmatrix}(-1)^{n-j-1-l}q_{2}^{c(n-j-1-l)}q_{4}^{c(l)}\right.$
@@
 $+$ $\left.\left.\eta^{\alpha c_{n}}\Bigl{(}\gamma(s\alpha^{\prime},k-l+n-1)\gamma(u\alpha^{\prime},-k+l+n-1)q_{2}^{a_{n}}\right.\right.$
--- $\left.\left.\gamma(s\alpha^{\prime},k-l+n)\gamma(u\alpha^{\prime},-k+l+n-2)q_{4}^{a_{n}}\Bigr{)}\right]\right\}$ (59)
+-- $\left.\left.\gamma(s\alpha^{\prime},k-l+n)\gamma(u\alpha^{\prime},-k+l+n-2)q_{4}^{a_{n}}\Bigr{)}\right]\right\}$
 and
-${\cal A}^{\alpha\beta(4)}_{4}({\bf 1},2,{\bf 3},4)$ $=$ $2i\alpha^{\prime}e^{2}(n!)^{2}\sum_{j=0}^{n}\frac{(2\alpha^{\prime})^{n-j}}{j![(n-j)!]^{2}}\eta^{a(j)c(j)}$
+${\cal A}^{\alpha\beta(4)}_{4}({\bf 1},2,{\bf 3},4)$ $=$ $2i\alpha^{\prime}e^{2}(n!)^{2}\sum_{j=0}^{n}\frac{(2\alpha^{\prime})^{n-j}}{j![(n-j)!]^{2}}\eta^{a(j)c(j)}$ (61)
 $\times$ $\left\{-\sum_{k=0}^{n-j}\begin{pmatrix}n-j\\ k\end{pmatrix}(-1)^{k}q_{2}^{a(n-j-k)}q_{4}^{a(k)}\sum_{l=0}^{n-j}\begin{pmatrix}n-j\\ l\end{pmatrix}(-1)^{n-j-l}q_{2}^{c(n-j-l)}q_{4}^{c(l)}\right.$
@@
 $\times$ $\left.\gamma(s\alpha^{\prime},k-l+n-1)\gamma(u\alpha^{\prime},-k+l+n-1)\biggr{]}\right.$
--- $\left.\frac{(n-j)}{(2\alpha^{\prime})}\sum_{k=0}^{n-j-1}\begin{pmatrix}n-j-1\\ k\end{pmatrix}(-1)^{k}q_{2}^{a(n-j-1-k)}q_{4}^{a(k)}\right.$ (67)
+-- $\left.\frac{(n-j)}{(2\alpha^{\prime})}\sum_{k=0}^{n-j-1}\begin{pmatrix}n-j-1\\ k\end{pmatrix}(-1)^{k}q_{2}^{a(n-j-1-k)}q_{4}^{a(k)}\right.$
 $\times$ $\left.\sum_{l=0}^{n-j-1}\begin{pmatrix}n-j-1\\ l\end{pmatrix}(-1)^{n-j-1-l}q_{2}^{c(n-j-1-l)}q_{4}^{c(l)}\right.$
@@
 -- $\left.\left.\gamma(s\alpha^{\prime},k-l+n)\gamma(u\alpha^{\prime},-k+l+n-2)q_{4}^{a_{n}}\right)\biggr{)}\biggr{]}\right.$
-$+$ $\left.\frac{(n-j)}{(2\alpha^{\prime})}\sum_{k=0}^{n-j-1}\begin{pmatrix}n-j-1\\ k\end{pmatrix}(-1)^{k}q_{2}^{a(n-j-1-k)}q_{4}^{a(k)}\right.$ (73)
+$+$ $\left.\frac{(n-j)}{(2\alpha^{\prime})}\sum_{k=0}^{n-j-1}\begin{pmatrix}n-j-1\\ k\end{pmatrix}(-1)^{k}q_{2}^{a(n-j-1-k)}q_{4}^{a(k)}\right.$
 $\times$ $\left.\sum_{l=0}^{n-j-1}\begin{pmatrix}n-j-1\\ l\end{pmatrix}(-1)^{n-j-1-l}q_{2}^{c(n-j-1-l)}q_{4}^{c(l)}\right.$
@@
 -- $\left.\left.\gamma(s\alpha^{\prime},k-l+n-1)\gamma(u\alpha^{\prime},-k+l+n)q_{4}^{a_{n}}\right)\Bigr{)}\biggr{]}\right.$
-$+$ $\left.\frac{(n-j)(n-j-1)\alpha^{\prime}t}{(2\alpha^{\prime})^{2}}\sum_{k=0}^{n-j-2}\begin{pmatrix}n-j-2\\ k\end{pmatrix}(-1)^{k}q_{2}^{a(n-j-2-k)}q_{4}^{a(k)}\right.$ (79)
+$+$ $\left.\frac{(n-j)(n-j-1)\alpha^{\prime}t}{(2\alpha^{\prime})^{2}}\sum_{k=0}^{n-j-2}\begin{pmatrix}n-j-2\\ k\end{pmatrix}(-1)^{k}q_{2}^{a(n-j-2-k)}q_{4}^{a(k)}\right.$
 $\times$ $\left.\sum_{l=0}^{n-j-2}\begin{pmatrix}n-j-2\\ l\end{pmatrix}(-1)^{n-j-2-l}q_{2}^{c(n-j-2-l)}q_{4}^{c(l)}\right.$
@@
 $+$ $\left.\gamma(s\alpha^{\prime},k-l+n)\gamma(u\alpha^{\prime},-k+l+n-2)q_{4}^{a_{n-1}}q_{4}^{a_{n}}\Bigr{)}\biggr{]}\right.$
-$+$ $\left.\frac{(n-j)^{2}\alpha^{\prime}t}{(2\alpha^{\prime})^{2}}\sum_{k=0}^{n-j-1}\begin{pmatrix}n-j-1\\ k\end{pmatrix}(-1)^{k}q_{2}^{a(n-j-1-k)}q_{4}^{a(k)}\right.$ (85)
+$+$ $\left.\frac{(n-j)^{2}\alpha^{\prime}t}{(2\alpha^{\prime})^{2}}\sum_{k=0}^{n-j-1}\begin{pmatrix}n-j-1\\ k\end{pmatrix}(-1)^{k}q_{2}^{a(n-j-1-k)}q_{4}^{a(k)}\right.$
 $\times$ $\left.\sum_{l=0}^{n-j-1}\begin{pmatrix}n-j-1\\ l\end{pmatrix}(-1)^{n-j-1-l}q_{2}^{c(n-j-1-l)}q_{4}^{c(l)}\right.$
@@
 $\times$ $\left.(-s\alpha^{\prime}+k-l+n)(-u\alpha^{\prime}-k+l+n)\right.$
-$\times$ $\left.\frac{1}{(\alpha^{\prime}t+1)}\gamma(s\alpha^{\prime},k-l+n-1)\gamma(u\alpha^{\prime},-k+l+n-1)\right\}\eta^{a_{S}c_{S}}$ (94)
+$\times$ $\left.\frac{1}{(\alpha^{\prime}t+1)}\gamma(s\alpha^{\prime},k-l+n-1)\gamma(u\alpha^{\prime},-k+l+n-1)\right\}\eta^{a_{S}c_{S}}$
 where $\gamma(x,N)\equiv\prod_{m=1}^{N}(-x+m)=(1-x)_{N}$, $q^{a(m)}\equiv q^{a_{1}}q^{a_{2}}\cdots q^{a_{m}}$, etc., and also we defined $\eta^{a(j)c(j)}\equiv\eta^{a_{1}c_{1}}\eta^{a_{2}c_{2}}\cdots\eta^{a_{j}c_{j}}$.
@@
 and so there are no tachyons in the spectrum, as expected.
+Moreover, the coupling constant $e$ arises by demanding that the amplitude describes the scattering between massive charged particles and photons – this requirement produces $e^{-\lambda}=\frac{1}{e^{2}}$.
 The other color-ordered amplitudes are given by
@@
 Finally, the momentum kernel ${\cal S}_{\alpha^{\prime}}$ appearing in the KLT formula given in the main text is given by [150]
-${\cal S}_{\alpha^{\prime}}[i_{1},\ldots,i_{k}|j_{1},\ldots j_{k}]_{p}\equiv\left(\frac{\pi\alpha^{\prime}}{2}\right)^{-k}\prod_{t=1}^{k}\sin\left[\frac{\pi\alpha^{\prime}}{2}\left(p\cdot k_{i_{t}}+\sum_{q>t}^{k}\theta(i_{t},i_{q})k_{i_{t}}\cdot k_{i_{q}}\right)\right]$ (101)
+${\cal S}_{\alpha^{\prime}}[i_{1},\ldots,i_{k}|j_{1},\ldots j_{k}]_{p}\equiv\left(\frac{\pi\alpha^{\prime}}{4}\right)^{-k}\prod_{t=1}^{k}\sin\left[\frac{\pi\alpha^{\prime}}{2}\left(p\cdot k_{i_{t}}+\sum_{q>t}^{k}\theta(i_{t},i_{q})k_{i_{t}}\cdot k_{i_{q}}\right)\right]$ (68)
 where $\theta(i_{t},i_{q})$ equals 1 if the ordering of the legs $i_{t}$ and $i_{q}$ is opposite in the sets $\{i_{1},\ldots,i_{k}\}$ and $\{j_{1},\ldots,j_{k}\}$, and zero if the ordering is the same.